### v2

# Reweight-annealing method for calculating the value of partition function via quantum Monte Carlo

Abstract
Efficient and accurate algorithm for partition function, free energy and thermal entropy calculations is of great significance in statistical physics and quantum many-body physics. Here we present an unbiased but low-technical-barrier algorithm within the quantum Monte Carlo framework, which has exceptionally high accuracy and no systemic error. Compared with the conventional specific heat integral method and Wang-Landau sampling algorithm, our method can obtain a much more accurate result of the sub-leading coefficient of the entropy consistent with CFT. This method can be widely used in both classical and quantum Monte Carlo simulations and is easy to be parallelized on computer.
[…]
On the other hand, the high-precision PF and its related quantities, such as free energy and entropy, are significantly important in numerical calculations since it means more effective information. For phase transitions no matter within or beyond the Ginzburg-Landau paradigm, an efficient and accurate way to straightforwardly obtain FP would be crucial, because its (high order) derivative reveals potential phase transitions. Besides, the precision of entropy can even reveal important intrinsic physics. In conformal field theory (CFT), the prefactor in front of the sub-leading term of entropy often carries universal information Calabrese and Lefevre (2008); Fradkin and Moore (2006); Nussinov and Ortiz (2009a, b); Casini and Huerta (2007); Ji and Wen (2019, 2020); Kong et al. (2020); Wu et al. (2021a); Zhao et al. (2021); Wu et al. (2021b). The thermal entropy also relates to the distinctiveness of many novel condensed matters, such as spin ice Snyder et al. (2001); Skjærvø et al. (2020); Udagawa et al. (2021); Bramwell and Harris (2020); Udagawa et al. (2021), topological order Jiang et al. (2012); Castelnovo et al. (2012); Sachdev (2018); Levin and Wen (2006); Yan et al. (2021, 2022), fracton phases Pretko et al. (2020); Nandkishore and Hermele (2019); Gromov and Radzihovsky (2024); Zhou et al. (2022), etc. Even a small numerical error may make our concerned quantities deviate a lot because the sub-leading term is much more sensitive than the leading one.
In this paper, we present a more accurate and efficient method for calculating PF, dubbed reweight-annealing algorithm, within the framework of Monte Carlo (MC) simulation, either quantum or classical. It is inspired by thermal and quantum annealing algorithm in the quantum simulation area Huse and Fisher (1986); Santoro et al. (2002); Lucas (2014); Kadowaki and Nishimori (1998); Santoro et al. (2002); Mézard and Montanari (2009); Heim et al. (2015); Yan et al. (2023); Ding et al. (2023) and recent developments in the entanglement calculations of QMC Alba (2017); D’Emidio (2020); Zhao et al. (2022a, b); Pan et al. (2023); Zhou et al. (2024). Our method not only has high accuracy for calculating PF, but is unbiased, easy to implement and naturally can be parallelized on computer.
Basic idea.- Given a partition function $Z(\beta)$, where $\beta$ is a general parameter which is not limited to temperature, but can be some other parameter in the Hamiltonian. For two parameter points $\beta^{\prime}$ and $\beta^{\prime\prime}$, we define the ratio
$\gamma(Z;\beta^{\prime},\beta^{\prime\prime})\equiv\frac{Z(\beta^{\prime\prime})}{Z(\beta^{\prime})}=\bigg{\langle}\frac{W(\beta^{\prime\prime})}{W(\beta^{\prime})}\bigg{\rangle}_{MC}$ (1)
where $W(\beta^{\prime})$ and $W(\beta^{\prime\prime})$ denote the weight of the same configuration sampled by MC simulation but with different parameters $\beta^{\prime}$ and $\beta^{\prime\prime}$. In this way, the ratio of partition functions with different parameters can be simulated via Monte Carlo in principle.
Figure 1: Schematic diagram of the reweight-annealing method. (a) Key point 1: Find an annealing path in the whole parameter space to connect the unknown point we want to solved and a known point. $T$ means temperature and $s$ is a general parameter of the Hamiltonian. (b) Key point 2: Use a sampled distribution (red) to represent another distribution (blue) through reweighting a same configuration. If these two distributions are close to each other, the effect of the reweighting would be good since the importance sampling can approximately be kept.
For the first consideration, we take $\beta$ as the inverse temperature. Notice that $Z(0)=d^{N}$ is pre-knowledge because all the configurations are equally important at infinite high temperature, where $N$ and $d$ are the number of particles and the freedom of a single particle respectively. Then we can obtain any $Z(\beta^{\prime})$ from Eq. (1) by setting $\beta^{\prime\prime}=0$. Therefore, the first key point of the method is to choose a known point $Z(\beta^{\prime\prime})$ as the reference which connects to some unknown point $Z(\beta^{\prime})$, as shown in Fig. 1 (a). The parameter of the path is not limited to the temperature $T=1/\beta$, once the parameter path can connect to a known point in the whole parameter space, then the method would work.
The second key point is keeping the importance sampling. Obviously, if the $\beta^{\prime}$ and $\beta^{\prime\prime}$ are close enough, the ratio of PF is close to 1 and is easy for calculation. Otherwise, the ratio would be close to zero or infinite which can not be well-sampled by MC. It can be easily understood in the reweighting frame as shown in Fig. 1 (b), if we want to use a well-known distribution $Z(\beta^{\prime})=\sum W(\beta^{\prime})$ to calculate another distribution $Z(\beta^{\prime\prime})=\sum W(\beta^{\prime\prime})$ via resetting the weight of samplings, the weight before/after resetting of the same configuration should be close to each other. It is still an importance sampling in this sense and requires the $\beta^{\prime}$ and $\beta^{\prime\prime}$ be close enough. As we will discuss later, the practical calculation of Eq. (1) should be divided into many steps for better accuracy, and the inverse temperature $\beta$ will slowly decrease from our concerned point $\beta^{\prime}$ to zero, which has some analogousness with the thermal or quantum annealing algorithm.
When we are considering the vicinity of the zero temperature, the inverse temperature span would be quite wide hence require a long runtime on computer. Thus we can find another alternative parameter path for the calculation according to the first key point mentioned above. One choice in this case is to fix the temperature as a constant. Instead, we introduce an extra term, e.g., transverse field $H_{1}=\sum_{i}\sigma_{i}^{x}$ to the target system $H_{0}$, and consider the partition function $Z^{\prime}(s)$ for $H(s)=sH_{0}+(1-s)H_{1}$, which is exactly the archetypal Hamiltonian of quantum annealing. Similarly, we define ${Z^{\prime}(s^{\prime\prime})}/{Z^{\prime}(s^{\prime})}=\langle{W(s^{\prime\prime})}/{W(s^{\prime})}\rangle_{MC}$ as our observable. The introduction of an external field is conducive because $H(0)$ only has one ground state, thus the entropy $S$ is minimized to be zero and $\ln Z^{\prime}(0)=-\lim_{\beta\to\infty}\beta E$, where $E=-N$ is the ground state energy of $H_{1}$.
Algorithm details.- In this paper, without loss of generality, we focus on the temperature case $\beta=1/T$, the counterpart of the thermal annealing, for demonstrations. We use the stochastic series expansion (SSE) in the followings, in which the weight ratio of temperature is ${W(\beta^{\prime\prime})}/{W(\beta^{\prime})}=({\beta^{\prime\prime}}/{\beta^{\prime}})^{n}$, where $n$ is the the order of series in the expansion. Since the value of the exact zero in the ratio is meaningless, practically, we set $\beta^{\prime\prime}$ in Eq. (1) a sufficiently small number $\beta_{0}$. The adequacy can be checked by comparing the results with that of a smaller $\beta_{0}$ so as to see whether the results ($\ln Z$) converge. Since the estimator in the right hand side of Eq. (1) is unbiased, the algorithm itself is unbiased. Attention that too small ratio of $(\beta_{0}/\beta^{\prime})^{n}$ not only requires more Monte Carlo time $N_{\text{MC}}$ since the error is proportional to $1/\sqrt{N_{\text{MC}}}$, but will exceed the data range of computer (e.g. $1.7\times 10^{-308}\sim 1.7\times 10^{308}$ for double type data).
To fix this problem, according to the second key point mentioned above, we divide $[\beta_{0},\beta^{\prime}]$ into $m$ segments, i.e.
$\frac{Z(\beta_{0})}{Z(\beta^{\prime})}=\prod_{k=1}^{m}\frac{Z(\beta_{k-1})}{Z(\beta_{k})}=\prod_{k=1}^{m}\gamma(Z;\beta_{k},\beta_{k-1})$ (2)
where $\beta_{m}\equiv\beta^{\prime}$. We evaluate each $\gamma(Z;\beta_{k},\beta_{k-1})$, and further group them into some batches. It also obviously reveals the method is naturally suitable for parallel computing. The value of each batch is the multiplication of all $\gamma(Z;\beta_{k},\beta_{k-1})$ in it, which we require to be in the same order of magnitude (e.g. $10^{-10}$ in our paper). This can avoid the loss of precision because finally we have to add all the logarithmic values of these batches.
For the multiplication in each batch, to ensure the preciseness, we define the schedule parameter $\alpha_{k}\equiv\beta_{k}/\beta_{k-1}>1$, and require
$\gamma(Z;\beta_{k},\beta_{k-1})=(\frac{1}{\alpha_{k}})^{n}\sim\epsilon$ (3)
According to the computation complexity of SSE a straightforward choice is to set the order of magnitude of $n\sim\mathcal{O}(\beta_{k}L)$ for typical 1D systems and $\mathcal{O}(\beta_{k}L^{2})$ for 2D systems. We can select an empirical and appropriate prefactor $\xi$, setting $n=\xi\beta_{k}L$ for 1D systems, then we do not need to determine $n$ from the thermalization part of Monte Carlo simulation, and the calculations for different $\gamma(Z;\beta_{k},\beta_{k-1})$ can be parallelized. The $\beta_{k}$ here plays similar role as the schedule parameter (temperature or the strength of magnetic field) in annealing algorithms which should vary slow enough to ensure ‘adiabaticity’ and keep the condition of importance sampling. If $\beta_{k}$ varies too dramatically, the accuracy of result will be influenced.
Attention that $n\geq 1$ in Eq. (3), therefore $\alpha_{k,\max}=1/\epsilon$. For low temperature points and large systems, $\alpha_{k}$ can be very close to $1$. To keep more significant digits, a lower bound for $\alpha_{k}$ is also needed, e.g., $\alpha_{k,\min}=1.00001$ in our simulations.
Figure 2: For $\beta=30$, the values of $\ln Z$ as a function of the chain length $L$. Our reweight-annealing (Re-An) method has perfect match with ED. For example, the ED and our results are 251.62283 and 251.624(2) respectively for $L=12$.
1D spin-1/2 Heisenberg chain.- As the first example, we consider a 1D spin-1/2 antiferromagnetic (AFM) Heisenberg chain with periodic boundary condition (PBC) at inverse temperature $\beta=30$. We set the $\epsilon=10^{-2}$ in Eq. 3. For the system size we simulated, the $\beta$ is large enough for checking the effectiveness of our method in low temperature. As it is shown in Fig. 2, results from our method can match that of the exact diagonalization (ED) method exceptionally well for small lengths in a very high-precision. The error bar is about 0.002 while the value is 251.624, and the relative error is smaller than $10^{-7}$. The result is also consistent with the expectation that $\ln Z\propto L+o(L)$, where $o(L)$ means higher order terms which is negligible compared with $L$. One may feel this example too trivial and easy to be calculated. Therefore we introduce the second example which needs a high-precision to extract the correction of the sub-leading term in entropy.
Figure 3: (a) Thermodynamic entropy $S(L)$ as a function of $L$ from different algorithms; (b) The $\ln L$ term extracted from calculating $2S(L)-S(2L)$, and the fitting slope (dashed green line) is 1.00(2).
2D LBW-HM.- To further exploit the advantage of our method, we calculate the thermodynamic entropy of a 2D lattice Bisognano-Wichmann Heisenberg model (LBW-HM) on a square cylinder Giudici et al. (2018); Dalmonte et al. (2018). The thermodynamic entropy $S$ of the LBW-HM at $\beta=1$ provides a good approximation for the Von Neumann entanglement entropy of a 2D PBC square lattice spin-1/2 AFM Heisenberg model under smooth bipartition, i.e., the concerned subsystem is a cylinder Mendes-Santos et al. (2019, 2020).
According to the CFT prediction and previous numerical results Metlitski and Grover (2015); Laflorencie (2016); Zhao et al. (2022a); Deng et al. (2023), the coefficient of sub-leading term of the entropy reflects the number of Goldstone modes. In this example, coarse result of entropy is disfavored because the sub-leading term in entropy is sensitive to errors. For this kind of problems, our method has great advantage over the method of specific heat integral ($C$-int) and Wang-Landau (WL) algorithm Wang and Landau (2001); Troyer et al. (2003) due to its high accuracy and efficiency.
We consider the same LBW-HM as Ref. Mendes-Santos et al. (2020); Giudici et al. (2018) which highly resembles the entanglement Hamiltonian of a spin-1/2 AFM Heisenberg model on a torus with cornerless cutting. Theoretical investigations predict the thermal entropy of the LBW-HM at $\beta=1$ obeys $S(L)=aL+b\ln L+c$, and $b=n_{G}/2$, where $n_{G}=2$ is the number of Goldstone modes for the 2D spin-1/2 AFM Heisenberg model Metlitski and Grover (2011); Laflorencie (2016). Therefore, we can extract $b$ from the slope of equation $2S(L)-S(2L)$, in which the leading term of $L$ can be cancelled with each other.
In Fig. 3, we compare the performance of different algorithms. Generally, the accuracy of the $C$-int method depends on how many intervals we divide. To compare with our method, we require the division of $\beta$ in the $C$-int method to be exactly the same with our method. As it is shown in Fig. 3, this introduces great systemic error because the division number in our method is not adequate for $C$-int. Therefore, to achieve a better result via $C$-int, more computation resources are needed compared to our method.
The WL algorithm, on another hand, suffers from the problem of error saturation, which though can be alleviated by some means like the $t^{-1}$ modification, where $t$ is the Monte Carlo time, the cut-off of the series expansion and the initial value of the refinement parameter $f$ are both hyperparameters to be determined Belardinelli et al. (2008); Belardinelli and Pereyra (2007); Zhou and Bhatt (2005); Inglis and Melko (2013). Therefore, it is very difficult to achieve results with an arbitrary precision with this method and the convergence time is not generally well-controlled. Besides, in Wang-Landau sampling, the change of $f$ typically obeys some monotonically decreasing function, and the computation subroutines with different $f$’s cannot be parallelized because of the linear logic of the algorithm. In this paper, we consider the vanilla WL algorithm, following Ref. Troyer et al. (2003), as an example. Fig. 3(a) shows that the WL algorithm gives roughly correct result for small sizes while the extraction of entropy in Fig. 3(b) completely reveals its unpreciseness, and the $b$ fitted from it is far from the theoretical prediction.
2D XXZ model.- As the last example, we consider a much more complex model, a 2D spin-1/2 XXZ model with next nearest neighbor interaction,
$H=-\sum_{\langle i,j\rangle}[2S^{z}_{i}S^{z}_{j}+S^{x}_{i}S^{x}_{j}+S^{y}_{i}S^{y}_{j}]$
$-0.2\sum_{\langle\langle i,j\rangle\rangle}[2S^{z}_{i}S^{z}_{j}+S^{x}_{i}S^{x}_{j}+S^{y}_{i}S^{y}_{j}].$ (4)
where $\langle i,j\rangle$ and $\langle\langle i,j\rangle\rangle$ denotes the nearest and the next nearest neighbors respectively. The lattice we simulated is a square lattice with PBC. This model has a finite temperature 2D Ising phase transition which can be captured by the second or higher order derivative of the free energy $F=-T\ln Z$.
Figure 4: (a) The second order derivative of the free energy, $\partial^{2}F/\partial T^{2}$, changes with the temperature $T$ in different system size. The peak diverges more when the size becomes larger, which probes the phase transition point here; (b) The binder ratios in this model at different sizes intersecting at a point which is highly consistent with the peak obtained by Re-An algorithm.
For convenience, here we use equal distance of $T=1/\beta=0.01$ to implement our Re-An method, and it is also easy for numerical differentiation. Because a constant factor of the PF does not affect the (higher order) derivative of $F$, thus we choose $Z(T^{*}=1.95)$ as the reference to replace annealing to a known point. The PF we gained is actually $Z(T)/Z(T^{*})$ where $Z(T^{*})$ is an unknown constant. Then we indeed observe a divergent peak of the second order derivative of the free energy as shown in Fig. 4(a), which accurately probes the phase transition point, matching with the intersection point of binder ratios $R_{2}=\langle m^{4}\rangle/\langle m^{2}\rangle^{2}$ of different sizes shown in Fig. 4(b).
This example also inspires us that connecting to a known point is not necessary when we just need the second or higher order derivative of free energy. The data also demonstrate the high-precision of our method, even the second order derivative is still smooth and precise.
Summary.- We propose an unbiased algorithm within Monte Carlo framework, which has low-technical barrier and no systemic error. The method can be easily implemented for parallel computation and can obtain the value of partition function, free energy and thermal entropy with high precision. The method has two keynotes: 1. Find a parameter path in whole space (either temperature or external parameter) connecting a known point and the to-be-solved one. If we only want to calculate the (high order) derivative of the PF, even fixing an unknown point as the reference is enough; 2. Annealing along the path to calculate the final result through a propagation formula.
We show the advantage of our method via calculating the thermal entropy of 2D LBW-HM and successfully extract precise coefficient of the sub-leading term ($\ln L$), which is consistent with the CFT prediction and catch the correct number of Goldstone modes. In comparison, the $C$-int and the WL sampling both have worse performance. In a 2D XXZ model, the numerical second derivative of the Re-An data is still smooth and precise, which accurately probes the phase transition point. Our method make the efficient and high-precision calculation of entropy and other related physical quantities accessible, which can reveal some universal information, such as CFT, criticality and symmetry breaking.
Acknowledgement.- We acknowledge the start-up funding of Westlake University. The authors also acknowledge Beijing PARATERA Tech Co.,Ltd.(https://www.paratera.com/) for providing HPC resources that have contributed to the research results reported within this paper. CC and ZY thank the support of the National Natural Science Foundation of China (grant nos. 12174167, 12247101).
Note: After we finished this work, we found a related work posted on arXiv two weeks ago Dai and Xu (2024).

### v3

@@
 On the other hand, the high-precision PF and its related quantities, such as free energy and entropy, are significantly important in numerical calculations since it means more effective information.
-For phase transitions no matter within or beyond the Ginzburg-Landau paradigm, an efficient and accurate way to straightforwardly obtain FP would be crucial, because its (high order) derivative reveals potential phase transitions.
-Besides, the precision of entropy can even reveal important intrinsic physics.
-In conformal field theory (CFT), the prefactor in front of the sub-leading term of entropy often carries universal information Calabrese and Lefevre (2008); Fradkin and Moore (2006); Nussinov and Ortiz (2009a, b); Casini and Huerta (2007); Ji and Wen (2019, 2020); Kong et al.
-(2020); Wu et al.
-(2021a); Zhao et al.
-(2021); Wu et al.
-(2021b).
+For phase transitions no matter within or beyond the Ginzburg-Landau paradigm, an efficient and accurate way to straightforwardly obtain PF would be crucial, because its (high order) derivative reveals potential phase transitions.
 The thermal entropy also relates to the distinctiveness of many novel condensed matters, such as spin ice Snyder et al.
@@
 (2022), etc.
+Besides, the precision of entropy can even reveal important intrinsic physics.
+In conformal field theory (CFT), the prefactor in front of the sub-leading term of entropy often carries universal information Calabrese and Lefevre (2008); Fradkin and Moore (2006); Nussinov and Ortiz (2009a, b); Casini and Huerta (2007); Ji and Wen (2019, 2020); Kong et al.
+(2020); Wu et al.
+(2021a); Zhao et al.
+(2021); Wu et al.
+(2021b).
 Even a small numerical error may make our concerned quantities deviate a lot because the sub-leading term is much more sensitive than the leading one.
@@
 (2023); Ding et al.
-(2023) and recent developments in the entanglement calculations of QMC Alba (2017); D’Emidio (2020); Zhao et al.
+(2023) and recent developments of the incremental trick in the entanglement calculations of QMC Alba (2017); D’Emidio (2020); Zhao et al.
 (2022a, b); Pan et al.
@@
 Our method not only has high accuracy for calculating PF, but is unbiased, easy to implement and naturally can be parallelized on computer.
-Basic idea.- Given a partition function $Z(\beta)$, where $\beta$ is a general parameter which is not limited to temperature, but can be some other parameter in the Hamiltonian.
-For two parameter points $\beta^{\prime}$ and $\beta^{\prime\prime}$, we define the ratio
-$\gamma(Z;\beta^{\prime},\beta^{\prime\prime})\equiv\frac{Z(\beta^{\prime\prime})}{Z(\beta^{\prime})}=\bigg{\langle}\frac{W(\beta^{\prime\prime})}{W(\beta^{\prime})}\bigg{\rangle}_{MC}$ (1)
-where $W(\beta^{\prime})$ and $W(\beta^{\prime\prime})$ denote the weight of the same configuration sampled by MC simulation but with different parameters $\beta^{\prime}$ and $\beta^{\prime\prime}$.
-In this way, the ratio of partition functions with different parameters can be simulated via Monte Carlo in principle.
+Basic idea.- Given a partition function $Z({\mathbf{p}})$, where ${\mathbf{p}}=[p_{1},p_{2},\cdots]$ denotes the parameters in the PF such as the temperature $T$ and some Hamiltonian parameter $s$.
+For example, Fig.1 (a) shows a diagram of the two parameters.
+For two different parameter points ${\mathbf{p}}^{\prime}$ and ${\mathbf{p}}^{\prime\prime}$, we define the reweighting ratio which can be measured in Monte Carlo
+$\frac{Z({\mathbf{p}}^{\prime})}{Z({\mathbf{p}}^{\prime\prime})}=\bigg{\langle}\frac{W({\mathbf{p}}^{\prime})}{W({\mathbf{p}}^{\prime\prime})}\bigg{\rangle}_{{\mathbf{p}}^{\prime\prime}}$ (1)
+where $W({\mathbf{p}}^{\prime})$ and $W({\mathbf{p}}^{\prime\prime})$ denote the weight of the same configuration under different parameters ${\mathbf{p}}^{\prime}$ and ${\mathbf{p}}^{\prime\prime}$, and the subscript indicates the result comes from the MC simulation under parameter ${\mathbf{p}}^{\prime\prime}$.
+As an example, we take ${\mathbf{p}}=\beta$, where $\beta$ denotes the inverse temperature.
+Under the framework of stochastic series expansion (SSE), we can then derive $Z(\beta^{\prime})/Z(\beta^{\prime\prime})=\langle(\beta^{\prime}/\beta^{\prime\prime})^{n}\rangle_{\beta^{\prime\prime}}$, where $n$ is the number of non-identity operators in the SSE simulation.
+The details of the $Z(\beta^{\prime})/Z(\beta^{\prime\prime})$ have been explained in the supplemental materials (SM).
 Figure 1: Schematic diagram of the reweight-annealing method.
-(a) Key point 1: Find an annealing path in the whole parameter space to connect the unknown point we want to solved and a known point.
-$T$ means temperature and $s$ is a general parameter of the Hamiltonian.
-(b) Key point 2: Use a sampled distribution (red) to represent another distribution (blue) through reweighting a same configuration.
+(a) Finding a reference point and the annealing path connecting the unknown point to it in the parameter space.
+$T$ and $s$ correspond to the thermal and quantum version, respectively; (b) Using a sampled distribution (red) to represent another distribution (blue) through reweighting a same configuration.
 If these two distributions are close to each other, the effect of the reweighting would be good since the importance sampling can approximately be kept.
-For the first consideration, we take $\beta$ as the inverse temperature.
-Notice that $Z(0)=d^{N}$ is pre-knowledge because all the configurations are equally important at infinite high temperature, where $N$ and $d$ are the number of particles and the freedom of a single particle respectively.
-Then we can obtain any $Z(\beta^{\prime})$ from Eq.
-(1) by setting $\beta^{\prime\prime}=0$.
-Therefore, the first key point of the method is to choose a known point $Z(\beta^{\prime\prime})$ as the reference which connects to some unknown point $Z(\beta^{\prime})$, as shown in Fig.
-1 (a).
-The parameter of the path is not limited to the temperature $T=1/\beta$, once the parameter path can connect to a known point in the whole parameter space, then the method would work.
-The second key point is keeping the importance sampling.
-Obviously, if the $\beta^{\prime}$ and $\beta^{\prime\prime}$ are close enough, the ratio of PF is close to 1 and is easy for calculation.
-Otherwise, the ratio would be close to zero or infinite which can not be well-sampled by MC.
-It can be easily understood in the reweighting frame as shown in Fig.
-1 (b), if we want to use a well-known distribution $Z(\beta^{\prime})=\sum W(\beta^{\prime})$ to calculate another distribution $Z(\beta^{\prime\prime})=\sum W(\beta^{\prime\prime})$ via resetting the weight of samplings, the weight before/after resetting of the same configuration should be close to each other.
-It is still an importance sampling in this sense and requires the $\beta^{\prime}$ and $\beta^{\prime\prime}$ be close enough.
-As we will discuss later, the practical calculation of Eq.
-(1) should be divided into many steps for better accuracy, and the inverse temperature $\beta$ will slowly decrease from our concerned point $\beta^{\prime}$ to zero, which has some analogousness with the thermal or quantum annealing algorithm.
-When we are considering the vicinity of the zero temperature, the inverse temperature span would be quite wide hence require a long runtime on computer.
-Thus we can find another alternative parameter path for the calculation according to the first key point mentioned above.
-One choice in this case is to fix the temperature as a constant.
-Instead, we introduce an extra term, e.g., transverse field $H_{1}=\sum_{i}\sigma_{i}^{x}$ to the target system $H_{0}$, and consider the partition function $Z^{\prime}(s)$ for $H(s)=sH_{0}+(1-s)H_{1}$, which is exactly the archetypal Hamiltonian of quantum annealing.
-Similarly, we define ${Z^{\prime}(s^{\prime\prime})}/{Z^{\prime}(s^{\prime})}=\langle{W(s^{\prime\prime})}/{W(s^{\prime})}\rangle_{MC}$ as our observable.
-The introduction of an external field is conducive because $H(0)$ only has one ground state, thus the entropy $S$ is minimized to be zero and $\ln Z^{\prime}(0)=-\lim_{\beta\to\infty}\beta E$, where $E=-N$ is the ground state energy of $H_{1}$.
-Algorithm details.- In this paper, without loss of generality, we focus on the temperature case $\beta=1/T$, the counterpart of the thermal annealing, for demonstrations.
-We use the stochastic series expansion (SSE) in the followings, in which the weight ratio of temperature is ${W(\beta^{\prime\prime})}/{W(\beta^{\prime})}=({\beta^{\prime\prime}}/{\beta^{\prime}})^{n}$, where $n$ is the the order of series in the expansion.
-Since the value of the exact zero in the ratio is meaningless, practically, we set $\beta^{\prime\prime}$ in Eq.
-(1) a sufficiently small number $\beta_{0}$.
-The adequacy can be checked by comparing the results with that of a smaller $\beta_{0}$ so as to see whether the results ($\ln Z$) converge.
-Since the estimator in the right hand side of Eq.
-(1) is unbiased, the algorithm itself is unbiased.
-Attention that too small ratio of $(\beta_{0}/\beta^{\prime})^{n}$ not only requires more Monte Carlo time $N_{\text{MC}}$ since the error is proportional to $1/\sqrt{N_{\text{MC}}}$, but will exceed the data range of computer (e.g.
-$1.7\times 10^{-308}\sim 1.7\times 10^{308}$ for double type data).
-To fix this problem, according to the second key point mentioned above, we divide $[\beta_{0},\beta^{\prime}]$ into $m$ segments, i.e.
-$\frac{Z(\beta_{0})}{Z(\beta^{\prime})}=\prod_{k=1}^{m}\frac{Z(\beta_{k-1})}{Z(\beta_{k})}=\prod_{k=1}^{m}\gamma(Z;\beta_{k},\beta_{k-1})$ (2)
-where $\beta_{m}\equiv\beta^{\prime}$.
-We evaluate each $\gamma(Z;\beta_{k},\beta_{k-1})$, and further group them into some batches.
-It also obviously reveals the method is naturally suitable for parallel computing.
-The value of each batch is the multiplication of all $\gamma(Z;\beta_{k},\beta_{k-1})$ in it, which we require to be in the same order of magnitude (e.g.
-$10^{-10}$ in our paper).
-This can avoid the loss of precision because finally we have to add all the logarithmic values of these batches.
-For the multiplication in each batch, to ensure the preciseness, we define the schedule parameter $\alpha_{k}\equiv\beta_{k}/\beta_{k-1}>1$, and require
-$\gamma(Z;\beta_{k},\beta_{k-1})=(\frac{1}{\alpha_{k}})^{n}\sim\epsilon$ (3)
-According to the computation complexity of SSE a straightforward choice is to set the order of magnitude of $n\sim\mathcal{O}(\beta_{k}L)$ for typical 1D systems and $\mathcal{O}(\beta_{k}L^{2})$ for 2D systems.
-We can select an empirical and appropriate prefactor $\xi$, setting $n=\xi\beta_{k}L$ for 1D systems, then we do not need to determine $n$ from the thermalization part of Monte Carlo simulation, and the calculations for different $\gamma(Z;\beta_{k},\beta_{k-1})$ can be parallelized.
-The $\beta_{k}$ here plays similar role as the schedule parameter (temperature or the strength of magnetic field) in annealing algorithms which should vary slow enough to ensure ‘adiabaticity’ and keep the condition of importance sampling.
-If $\beta_{k}$ varies too dramatically, the accuracy of result will be influenced.
-Attention that $n\geq 1$ in Eq.
-(3), therefore $\alpha_{k,\max}=1/\epsilon$.
-For low temperature points and large systems, $\alpha_{k}$ can be very close to $1$.
-To keep more significant digits, a lower bound for $\alpha_{k}$ is also needed, e.g., $\alpha_{k,\min}=1.00001$ in our simulations.
-Figure 2: For $\beta=30$, the values of $\ln Z$ as a function of the chain length $L$.
-Our reweight-annealing (Re-An) method has perfect match with ED.
-For example, the ED and our results are 251.62283 and 251.624(2) respectively for $L=12$.
-1D spin-1/2 Heisenberg chain.- As the first example, we consider a 1D spin-1/2 antiferromagnetic (AFM) Heisenberg chain with periodic boundary condition (PBC) at inverse temperature $\beta=30$.
-We set the $\epsilon=10^{-2}$ in Eq.
-3.
-For the system size we simulated, the $\beta$ is large enough for checking the effectiveness of our method in low temperature.
+The second step is to identify a reference point ${\mathbf{p}}_{0}$, of which the value of $Z({\mathbf{p}}_{0})$ has already been known.
+Therefore, by evaluating $Z({\mathbf{p}}_{0})/Z({\mathbf{p}}^{\prime\prime})$, we can directly determine the value of $Z({\mathbf{p}}^{\prime\prime})$.
+In the example that ${\mathbf{p}}=\beta$, apparently $Z(0)=d^{N}$ is pre-knowledge because all the configurations are equally important at infinite high temperature, where $N$ and $d$ are the number of particles and the degree of freedom of a single particle respectively.
+However, when ${\mathbf{p}}_{0}$ and ${\mathbf{p}}^{\prime\prime}$ are far away in the parameter space, the value of $Z({\mathbf{p}}_{0})/Z({\mathbf{p}}^{\prime\prime})$ would be close to zero or infinity, resulting in the problem of inefficiency and low precision for the MC simulation.
+This can be easily understood in the reweighting frame as shown in Fig.
+1 (b).
+If we want to use a well-known distribution $Z({\mathbf{p}}^{\prime\prime})=\sum W({\mathbf{p}}^{\prime\prime})$ to calculate another distribution $Z({\mathbf{p}}^{\prime})=\sum W({\mathbf{p}}^{\prime})$ via resetting the weight of samplings, the weight before/after the resetting of the same configuration should be close to each other.
+It is still an importance sampling in this sense and requires that ${\mathbf{p}}^{\prime}$ and ${\mathbf{p}}^{\prime\prime}$ should be close enough.
+Based on this consideration, practically, we divide the calculation of of Eq. (1) into many steps to enhance precision, i.e.
+$\frac{Z({\mathbf{p}}_{0})}{Z({\mathbf{p}}^{\prime\prime})}=\prod_{k=1}^{m}\frac{Z({\mathbf{p}}_{k-1})}{Z({\mathbf{p}}_{k})}$ (2)
+where ${\mathbf{p}}_{m}\equiv{\mathbf{p}}^{\prime\prime}$ and $Z({\mathbf{p}}_{k-1})/Z({\mathbf{p}_{k}})\to 1$ as long as ${\mathbf{p}}_{k}\to{\mathbf{p}}_{k-1}$.
+Here, ${\mathbf{p}}$ can be perceived to gradually change from the point of interest ${\mathbf{p}}^{\prime\prime}$ to the reference point ${\mathbf{p}}_{0}$ along a trajectory (annealing path) in the parameter space, akin to an adiabatic process or annealing as shown in Fig.1 (a).
+Therefore by calculating Eq.
+(2), not only we are able to achieve a higher precision of $Z({\mathbf{p}}^{\prime\prime})$, but concurrently track the variation of PF along this annealing path.
+This allows us to further analyze the variation of the free energy (and its derivatives) of the system.
+Apparently, the annealing path that ${\mathbf{p}}$ follows determines how large $m$ should we need and how ${\mathbf{p}}$ varies to perform the calculations of Eq.
+(2).
+It can be expected that, similar to conventional thermal annealing or quantum annealing, should the annealing trajectory approach a critical point, an increased number of divisions would be necessary in its vicinity.
+However, we point out that this actually affects little to our algorithm since numerical simulations only deal with finite system, where the criticality does not result in significant inefficiency.
+We have also verified this point in our simulations.
+For example, in the 2D spin-$1/2$ XXZ model which we present below, where ${\mathbf{p}}=T$ is the temperature, we implemented an equal division strategy with step size $\Delta T=0.01$.
+With $20000$ Monte Carlo samples for each bin, we are able to achieve a sufficiently smooth curve for further numerical differentiation.
+More discussions about the segmentation are in the section of “Dynamical annealing process”.
+Thermal and quantum reweight-annealing- As it shown in Fig.
+1, if ${\mathbf{p}}=\beta$, as previously mentioned, the annealing path connects an unknown inverse temperature point $\beta^{\prime\prime}$ to the infinite temperature point $\beta_{0}$, which serves as the reference.
+For convenience, we call this kind of process the thermal reweight-annealing (T-Re-An).
+Similarly, we can consider the so-called quantum reweight-annealing (Q-Re-An) by setting ${\mathbf{p}}=s$, where $s$ comes from $H(s)=sH_{0}+(1-s)H_{1}$, which is the archetypal Hamiltonian of quantum annealing with $Z(s)=e^{-\beta H(s)}$.
+Here, $H_{1}$ correspond to our reference point $s=0$, and we would like to obtain the PF related to $H_{0}$.
+We emphasis that different from the thermal version, the choice of $H_{1}$ here is much more flexible.
+For example, we can take $H_{1}=\sum_{i}\sigma_{i}^{x}$.
+Then at the zero temperature ($\beta\to\infty$), $H_{1}$ only has a unique ground state, thus its thermal entropy is minimized to be zero, resulting in $\ln Z(0)=-\lim_{\beta\to\infty}\beta E$, where $E=-N$ is the ground state energy of $H_{1}$.
+More details about the T-Re-An and Q-Re-An scheme can be found in the SM.
+Dynamical annealing process.- For the thermal and quantum reweight-annealing schemes, we provide a dynamical strategy to automatically regulate the alteration of ${\mathbf{p}}$ throughout the annealing process.
+This not only serves as an economic way for the division in Eq.
+(2), but can avoid the problem of precision loss when different $r({\mathbf{p}}_{k-1},{\mathbf{p}}_{k})$ differ by several orders of magnitude in the computer.
+For demonstration, we only discuss the thermal case, i.e.
+${\mathbf{p}}=\beta$, in this section.
+Without loss of generality, we assume $\beta_{k-1}<\beta_{k}$ in Eq. (2), and define the rate $\alpha_{k}\equiv\beta_{k-1}/\beta_{k}<1$. Within the framework of SSE, we have
+$\frac{Z(\beta_{k-1})}{Z(\beta_{k})}=\langle\alpha^{n_{k}}_{k}\rangle_{\beta_{k}}$ (3)
+where $n_{k}$ is the number of all non-identity operators in the SSE simulation, and the subscript $\beta_{k}$ denotes that this simulation is performed at the parameter point $\beta_{k}$.
+The annealing process starts from $k=m$, and we require $Z(\beta_{k-1})/Z(\beta_{k})=\epsilon<1$, where the choice of $\epsilon$ should match $N_{\text{MC}}$, the number of MC steps we use for samplings, as the error in MC simulations is proportional to $1/\sqrt{N_{\text{MC}}}$.
+Next, we set $n_{k}\approx\tilde{n}_{k}=\Lambda_{k}\beta_{k}L$ ($\Lambda_{k}\beta_{k}L^{2}$) for 1D (2D) system, and $\Lambda_{k}$ can be ascertained through the examination of result convergence.
+This comes from the fact that in SSE, the energy satisfies $E_{k}=-\langle n_{k}\rangle/\beta_{k}$.
+Practically, we set $\Lambda_{k}$ to be the same for all $k$ for convenience.
+This strategy not only makes $Z(\beta_{k-1})/Z(\beta_{k})\approx\epsilon$ for all $k$, but determines the next step in the annealing path, i.e.
+$\beta_{k-1}=\alpha_{k}\beta_{k}$.
+The division will be more sparse when $\beta$ gets smaller as we expected, where the thermal fluctuation has destroyed any order and $Z(\beta_{k-1})/Z(\beta_{k})\to 1$, which has little contribution to the result.
+Apparently, this strategy naturally allows for parallelization of evaluating different $Z(\beta_{k-1})/Z(\beta_{k})$ in Eq.
+(2).
+It is also easy to calculate that the number of divisions is around $(\beta-\beta_{0})\Lambda L/|\ln\epsilon|$ for large $L$ or $(\beta-\beta_{0})$, where $\Lambda_{k}=\Lambda$ for all $k$.
+Therefore, this dynamical annealing strategy does not require exponentially increasing number of divisions.
+Figure 2: For $\beta=30$, the values of $\ln Z$ as a function of the chain length $L$, where both T-Re-An method and Q-Re-An method have a perfect match with ED.
+For example, for $L=12$, the results from the two methods are 251.625(2) and 251.624(2), respectively, and the exact result is 251.62283 from ED calculations.
+Attention that here in practical SSE simulations, the Hamiltonian has a energy shift of $-1/4$ for each two-body term.
+1D spin-1/2 Heisenberg chain.- As the first example, we consider an 1D spin-1/2 antiferromagnetic (AFM) Heisenberg chain with periodic boundary condition (PBC) with $\beta=30$ and length $L$.
+We use both the T-Re-An method and the Q-Re-An method to calculate the PF.
+For the quantum version, we write
+$H(s)=s\bigg{[}\sum_{i}\mathbf{S}_{i}\cdot\mathbf{S}_{i+1}\bigg{]}+(1-s)\bigg{[}\sum_{i\in\text{even}}\mathbf{S}_{i}\cdot\mathbf{S}_{i+1}\bigg{]}$ (4)
+where $H_{0}=\sum_{i}\mathbf{S}_{i}\cdot\mathbf{S}_{i+1}$ and $H_{1}=\sum_{i\in\text{even}}{\mathbf{S}}_{i}\cdot{\mathbf{S}}_{i+1}$.
+Apparently, the unique ground state of $H_{1}$ is a tensor product of $L/2$ singlets, each carrying energy $-3/4$ ($\hbar=1$).
+The estimator for $Z(s^{\prime})/Z(s^{\prime\prime})$ is $\langle(s^{\prime}/s^{\prime\prime})^{n_{\text{odd}}}\rangle_{s^{\prime\prime}}$ in this model, with $n_{\text{odd}}$ denoting the amount of non-identity operators acting on all the bonds with odd indices within the framework of SSE.
 As it is shown in Fig.
-2, results from our method can match that of the exact diagonalization (ED) method exceptionally well for small lengths in a very high-precision.
-The error bar is about 0.002 while the value is 251.624, and the relative error is smaller than $10^{-7}$.
-The result is also consistent with the expectation that $\ln Z\propto L+o(L)$, where $o(L)$ means higher order terms which is negligible compared with $L$.
+2, both the results of T-Re-An and Q-Re-An method can match that of the exact diagonalization (ED) method exceptionally well for small lengths, which verify the validity of our methods.
+The results are also consistent with the expectation that $\ln Z\propto L+o(L)$, where $o(L)$ means higher order terms which is negligible compared with $L$.
 One may feel this example too trivial and easy to be calculated.
 Therefore we introduce the second example which needs a high-precision to extract the correction of the sub-leading term in entropy.
-Figure 3: (a) Thermodynamic entropy $S(L)$ as a function of $L$ from different algorithms; (b) The $\ln L$ term extracted from calculating $2S(L)-S(2L)$, and the fitting slope (dashed green line) is 1.00(2).
-2D LBW-HM.- To further exploit the advantage of our method, we calculate the thermodynamic entropy of a 2D lattice Bisognano-Wichmann Heisenberg model (LBW-HM) on a square cylinder Giudici et al.
+Figure 3: (a) Thermodynamic entropies $S(L)$ as a function of $L$ calculated from different algorithms; (b) The $\ln L$ term extracted by calculating $2S(L)-S(2L)$, and the fitting slope (dashed green line) here is 1.00(2).
+2D LBW-HM.- To further exploit the advantage of our method, we calculate the thermodynamic entropy of a 2D lattice Bisognano-Wichmann Heisenberg model (LBW-HM) Giudici et al.
 (2018); Dalmonte et al.
-(2018).
-The thermodynamic entropy $S$ of the LBW-HM at $\beta=1$ provides a good approximation for the Von Neumann entanglement entropy of a 2D PBC square lattice spin-1/2 AFM Heisenberg model under smooth bipartition, i.e., the concerned subsystem is a cylinder Mendes-Santos et al.
+(2018), the thermodynamic entropy $S$ of the which at $\beta=1$ provides a good approximation for the Von Neumann entanglement entropy of a 2D PBC square lattice spin-1/2 AFM Heisenberg model under smooth bipartition Mendes-Santos et al.
 (2019, 2020).
@@
 (2023), the coefficient of sub-leading term of the entropy reflects the number of Goldstone modes.
-In this example, coarse result of entropy is disfavored because the sub-leading term in entropy is sensitive to errors.
-For this kind of problems, our method has great advantage over the method of specific heat integral ($C$-int) and Wang-Landau (WL) algorithm Wang and Landau (2001); Troyer et al.
-(2003) due to its high accuracy and efficiency.
-We consider the same LBW-HM as Ref.
-Mendes-Santos et al.
-(2020); Giudici et al.
-(2018) which highly resembles the entanglement Hamiltonian of a spin-1/2 AFM Heisenberg model on a torus with cornerless cutting.
-Theoretical investigations predict the thermal entropy of the LBW-HM at $\beta=1$ obeys $S(L)=aL+b\ln L+c$, and $b=n_{G}/2$, where $n_{G}=2$ is the number of Goldstone modes for the 2D spin-1/2 AFM Heisenberg model Metlitski and Grover (2011); Laflorencie (2016).
-Therefore, we can extract $b$ from the slope of equation $2S(L)-S(2L)$, in which the leading term of $L$ can be cancelled with each other.
-In Fig.
-3, we compare the performance of different algorithms.
-Generally, the accuracy of the $C$-int method depends on how many intervals we divide.
-To compare with our method, we require the division of $\beta$ in the $C$-int method to be exactly the same with our method.
-As it is shown in Fig.
-3, this introduces great systemic error because the division number in our method is not adequate for $C$-int.
-Therefore, to achieve a better result via $C$-int, more computation resources are needed compared to our method.
-The WL algorithm, on another hand, suffers from the problem of error saturation, which though can be alleviated by some means like the $t^{-1}$ modification, where $t$ is the Monte Carlo time, the cut-off of the series expansion and the initial value of the refinement parameter $f$ are both hyperparameters to be determined Belardinelli et al.
-(2008); Belardinelli and Pereyra (2007); Zhou and Bhatt (2005); Inglis and Melko (2013).
-Therefore, it is very difficult to achieve results with an arbitrary precision with this method and the convergence time is not generally well-controlled.
-Besides, in Wang-Landau sampling, the change of $f$ typically obeys some monotonically decreasing function, and the computation subroutines with different $f$’s cannot be parallelized because of the linear logic of the algorithm.
-In this paper, we consider the vanilla WL algorithm, following Ref.
-Troyer et al.
-(2003), as an example.
+Here, the thermal entropy of the LBW-HM at $\beta=1$ should obey $S(L)=aL+b\ln L+c$, and $b=n_{G}/2$, where $n_{G}=2$ is the number of Goldstone modes of the 2D spin-1/2 AFM Heisenberg model Metlitski and Grover (2011); Laflorencie (2016).
+In this context, coarse result of entropy is disfavored because the the coefficient $b$ in front of the sub-leading term is sensitive to errors.
+We have compared our method with the specific heat integral ($C$-int) and Wang-Landau (WL) algorithm in this example.
+Similar to our method, the $C$-int method also depends on how many intervals we divide for $\beta$.
+Therefore we take the same division and Monte Carlo steps to compare the performance of these two algorithms.
+As it shown in Fig.
+3, the $C$-int method has great systemic error because the number of division is adequate for our method, but not the $C$-int method, indicating more computational resources needed for it to achieve better accuracy.
+For the WL algorithm, on another hand, the allocation of total computational resources should not be predetermined as it hinges on when we can achieve a converged and flat histogram.
 Fig.
-3(a) shows that the WL algorithm gives roughly correct result for small sizes while the extraction of entropy in Fig.
-3(b) completely reveals its unpreciseness, and the $b$ fitted from it is far from the theoretical prediction.
+3 shows great error of the results obtained from the vanilla WL samplings Troyer et al.
+(2003).
+This reflects the problem of error saturation, which is intrinsic in this algorithm due to an approximately flat histogram and the empirical refinement parameter Belardinelli et al.
+(2008); Belardinelli and Pereyra (2007); Zhou and Bhatt (2005).
+Another disadvantage of the WL sampling compared with our method is that it cannot be parallelized: the two subroutines for flattening the corresponding histograms before and after an adjustment of the refinement parameter, must be executed sequentially.
+To achieve an arbitrary precision within the Wang-Landau framework is not easy.
+In contrast, our method is much more accessible as we only need to measure the observable of $Z({\mathbf{p}}_{k-1})/Z({\mathbf{p}}_{k})$ in many conventional and mature MC algorithms.
 2D XXZ model.- As the last example, we consider a much more complex model, a 2D spin-1/2 XXZ model with next nearest neighbor interaction,
@@
 4(a), which accurately probes the phase transition point, matching with the intersection point of Binder ratios $R_{2}=\langle M^{4}\rangle/\langle M^{2}\rangle^{2}$ of different sizes shown in Fig.
-4(b).
+4(b), where $M$ denotes the magnetization.
 This example also inspires us that connecting to a known point is not necessary when we just need the second or higher order derivative of free energy.
-The data also demonstrate the high-precision of our method, even the second order derivative is still smooth and precise.
+The data also demonstrate the high-precision of our method, where the second order derivative is still smooth and precise.
 Summary.- We propose an unbiased algorithm within Monte Carlo framework, which has low-technical barrier and no systemic error.
 The method can be easily implemented for parallel computation and can obtain the value of partition function, free energy and thermal entropy with high precision.
-The method has two keynotes: 1.
-Find a parameter path in whole space (either temperature or external parameter) connecting a known point and the to-be-solved one.
-If we only want to calculate the (high order) derivative of the PF, even fixing an unknown point as the reference is enough; 2.
-Annealing along the path to calculate the final result through a propagation formula.
+The method has two keynotes: (1) Find a parameter path in whole space (either temperature or external parameter) connecting a known point and the to-be-solved one.
+If we only want to calculate the (high order) derivative of the PF, even fixing an unknown point as the reference is enough; (2) Annealing along the path to calculate the final result through a propagation formula.
 We show the advantage of our method via calculating the thermal entropy of 2D LBW-HM and successfully extract precise coefficient of the sub-leading term ($\ln L$), which is consistent with the CFT prediction and catch the correct number of Goldstone modes.
@@
 12174167, 12247101).
-Note: After we finished this work, we found a related work posted on arXiv two weeks ago Dai and Xu (2024).
+Appendix A Supplemental material
+A.1 (I) Estimators in thermal and quantum reweight-annealing under the SSE framework
+For the thermal case, the estimator of $Z(\beta^{\prime})/Z(\beta^{\prime\prime})$ is general, which is
+$\begin{split}\frac{Z(\beta^{\prime})}{Z(\beta^{\prime\prime})}=&\frac{1}{Z(\beta^{\prime\prime})}\text{tr}\{e^{-\beta^{\prime}H}\}\\ =&\frac{1}{Z(\beta^{\prime\prime})}\text{tr}\bigg{\{}\sum_{n}\bigg{(}\frac{\beta^{\prime}}{\beta^{\prime\prime}}\bigg{)}^{n}\frac{(\beta^{\prime\prime})^{n}}{n!}(-H)^{n}\bigg{\}}\\ =&\bigg{\langle}\bigg{(}\frac{\beta^{\prime}}{\beta^{\prime\prime}}\bigg{)}^{n}\bigg{\rangle}_{\beta^{\prime\prime}}\\ \end{split}$ (1)
+and the subscript denotes that this simulation should be performed at the inverse temperature $\beta^{\prime\prime}$.
+For the quantum case with $H(s)=sH_{0}+(1-s)H_{1}=H_{1}+s(H_{0}-H_{1})$, $Z(s^{\prime})/Z(s^{\prime\prime})$ usually takes the form $\langle(s^{\prime}/s^{\prime\prime})^{n_{s}}\rangle_{s^{\prime\prime}}$ within the framework of SSE, where $n_{s}$ is the number of non-identity operators related to $(H_{0}-H_{1})$.
+For instance, if
+$\begin{split}H(s)=&s\bigg{[}\sum_{i}\mathbf{S}_{i}\cdot\mathbf{S}_{i+1}\bigg{]}+(1-s)\bigg{[}\sum_{i\in\text{even}}\mathbf{S}_{i}\cdot\mathbf{S}_{i+1}\bigg{]}\\ =&\sum_{i\in\text{even}}\mathbf{S}_{i}\cdot\mathbf{S}_{i+1}+s\bigg{[}\sum_{i\in\text{odd}}\mathbf{S}_{i}\cdot\mathbf{S}_{i+1}\bigg{]}\end{split}$ (2)
+where $H_{0}=\sum_{i}\mathbf{S}_{i}\cdot\mathbf{S}_{i+1}$ and $H_{1}=\sum_{i\in\text{even}}{\mathbf{S}}_{i}\cdot{\mathbf{S}}_{i+1}$, then
+$\begin{split}\frac{Z(s^{\prime})}{Z(s^{\prime\prime})}=&\frac{1}{Z(s^{\prime\prime})}\text{tr}\{e^{-\beta H(s^{\prime})}\}\\ =&\frac{1}{Z(s^{\prime\prime})}\sum_{\alpha,S_{M}}\bigg{(}\frac{s^{\prime}}{2}\bigg{)}^{n_{\text{odd}}}\bigg{(}\frac{1}{2}\bigg{)}^{n_{\text{even}}}\tilde{S}_{\alpha,M}\\ =&\frac{1}{Z(s^{\prime\prime})}\sum_{\alpha,S_{M}}\bigg{(}\frac{s^{\prime}}{s^{\prime\prime}}\bigg{)}^{n_{\text{odd}}}\bigg{(}\frac{s^{\prime\prime}}{2}\bigg{)}^{n_{\text{odd}}}\bigg{(}\frac{1}{2}\bigg{)}^{n_{\text{even}}}\tilde{S}_{\alpha,M}\\ =&\bigg{\langle}\bigg{(}\frac{s^{\prime}}{s^{\prime\prime}}\bigg{)}^{n_{\text{odd}}}\bigg{\rangle}_{s^{\prime\prime}}\\ \end{split}$ (3)
+where $n_{\text{odd}}$ ($n_{\text{even}}$) denotes the amounts of non-identity operators acting on the bonds with odd (even) indices in SSE simulation, and
+$\tilde{S}_{\alpha,M}=\frac{\beta^{n}(M-n)!}{M!}\prod_{i=1}^{M}\langle\alpha_{i-1}|H_{a_{i},b_{i}}|\alpha_{i}\rangle$ (4)
+corresponds to operator string $S_{M}$ and the propagated states $\{\alpha_{0},\alpha_{1},\dots\alpha_{M}\}$ along the axis of imaginary time.
+A.2 (II) Technical details of the annealing process
+For the thermal and quantum reweight-annealing scheme with a scalar parameter $p$, our target is to compute
+$\frac{Z(p_{0})}{Z(p)}=\bigg{\langle}\frac{W(p_{0})}{W(p)}\bigg{\rangle}_{p}$ (5)
+where $p_{0}$ is the reference point in the parameter space.
+For the models we consider, $p_{0}=0$.
+Practically, we take $p_{0}$ to be a sufficiently small number (e.g.
+$10^{-8}$ in our paper) in the computer.
+The adequacy can be checked by comparing the results with that of a smaller $p_{0}$ so as to see whether the results converge.
+We first take $p=\beta$ for illustration and the case that $p=s$ is similar.
+The difference will be clarified at the end of this section.
+Without loss of generality, we assume $\beta_{0}<\beta$.
+In case that $\beta_{0}$ and $\beta$ are far away in the parameter space, to ensure importance sampling, we divide $[\beta_{0},\beta]$ into $m$ intervals $[\beta_{0},\beta_{1}]$, $[\beta_{1},\beta_{2}]$, $\cdots$, $[\beta_{m-1},\beta_{m}]$, where $\beta_{m}\equiv\beta$, i.e.
+$\frac{Z(\beta_{0})}{Z(\beta)}=\prod_{k=1}^{m}\frac{Z(\beta_{k-1})}{Z(\beta_{k})}$ (6)
+Each interval $[\beta_{k-1},\beta_{k}]$ ought to be sufficiently narrow so that the distribution at $\beta_{k}$ can well samples the configurations at $\beta_{k-1}$.
+Otherwise, the sampling process would be inefficient.
+Then the natural questions following are how to partition $[\beta_{0},\beta]$ and the how many intervals we need for Eq.
+(6)
+It can be expected that if $\beta_{k}$ and $\beta_{k-1}$ are separated on the two sides of the critical point, we may need a bulk of samples to estimate $Z(\beta_{k-1})/Z(\beta_{k})$.
+However, we point out that this actually affects little to our algorithm since numerical simulations only deal with finite system, where the criticality does not result in significant inefficiency.
+We have also verified this point in our simulations.
+For example, in the 2D spin-$1/2$ XXZ model, we implemented an equal division strategy with step size $\Delta T=0.01$.
+With $20000$ Monte Carlo samples for each bin, we were able to achieve a sufficiently smooth curve for further numerical differentiation.
+However, the strategy that we divide $[\beta_{0},\beta]$ into equal intervals or a geometric progression is not economic.
+We can expect that when $\beta\to 0$ (approaching infinite high temperature), the division can be more sparse compared to that at low temperatures.
+Therefore, an economic strategy to divide the interval is needed.
+In addition to the division problem, $Z(\beta_{0})/Z(\beta)$ could be an exponentially small number exceeding the data range of computer (e.g.
+$1.7\times 10^{-308}\sim 1.7\times 10^{308}$ for double type data), and we should store $\ln[Z(\beta_{0})/Z(\beta)]$ in the computer.
+However, simply applying the logarithm to Eq.
+(6) and converting the multiplication to addition may lead to precision loss in computer calculations if $\ln[Z(\beta_{k-1})/Z(\beta_{k})]$ and $\ln[Z(\beta_{k})/Z(\beta_{k+1})]s$ differ by several orders of magnitude.
+To address the two problems above, we present a dynamical approach here. Define $\alpha_{k}\equiv\beta_{k-1}/\beta_{k}<1$, and within the framework of SSE, we have
+$\frac{Z(\beta_{k-1})}{Z(\beta_{k})}=\langle\alpha^{n_{k}}_{k}\rangle_{\beta_{k}}$ (7)
+where $n_{k}$ is the number of all non-identity operators in the SSE simulation. Our dynamical process then starts from the last point with $k=m$, and we require
+$\frac{Z(\beta_{k-1})}{Z(\beta_{k})}=\epsilon<1,$ (8)
+where the choice of $\epsilon$ should match $N_{\text{MC}}$, the number of MC steps we use for samplings, as the error in MC simulations is proportional to $1/\sqrt{N_{\text{MC}}}$.
+Next, we set $n_{k}\approx\tilde{n}_{k}=\Lambda_{k}\beta_{k}L$ ($\Lambda_{k}\beta_{k}L^{2}$) for 1D (2D) system, and $\Lambda_{k}$ can be ascertained through the examination of result convergence.
+This comes from the fact that in SSE, the energy satisfies $E_{k}=-\langle n_{k}\rangle/\beta_{k}$.
+Practically, we set $\Lambda_{k}$ to be the same for all $k$ for convenience.
+Then $Z(\beta_{k-1})/Z(\beta_{k})$ can be obtained from the simulation from
+$\begin{split}\langle\alpha_{k}^{n_{k}}\rangle=\langle(\epsilon^{1/\tilde{n}_{k}})^{n_{k}}\rangle_{\beta_{k}}\approx\epsilon,\quad\beta_{k-1}=\alpha_{k}\beta_{k}\end{split}$ (9)
+Eq.
+(9) not only requires $Z(\beta_{k-1})/Z(\beta_{k})\approx\epsilon$ for all $k$, but determines the next step in the annealing path.
+Therefore, with this dynamical annealing strategy, the division will be more sparse when $\beta$ gets smaller.
+This strategy actually requires the number divisions of scale proportional to the system size for large systems. For the 1D system as an example,
+$\beta_{k-1}=\epsilon^{\frac{1}{\Lambda\beta_{k}L}}\beta_{k}\approx\bigg{(}1+\frac{\ln\epsilon}{\Lambda\beta_{k}L}\bigg{)}\beta_{k}$ (10)
+or
+$\beta_{k}-\beta_{k-1}=-\frac{\ln\epsilon}{\Lambda L}$ (11)
+then the number of divisions is around $(\beta-\beta_{0})\Lambda L/|\ln\epsilon|$.
+One more thing we have to pay attention is that when $\beta\to 0$, we have $\alpha_{k}\to 0$, which will also exceed the range of a number saved in the computer.
+Therefore we manually set a lower bound for $\alpha_{k}$ (e.g.
+$\min{\alpha_{k}}=\epsilon$ in our simulation).
+The existence of lower bound makes the division in the region that $\beta\to 0$) denser than that Eq.
+(9) requires.
+Since most $Z(\beta_{k-1})/Z(\beta_{k})$ still at the same order of magnitude, e.g.
+$\epsilon=10^{-2}$ used for the first two example in our simulation, and the more divisions when $\beta\to 0$ only constitutes a very small proportion, this affects little to the precision.
+In Fig.
+1, we plot the number of divisions in the T-Re-An method of first example in our paper for reference, where we take $\Lambda=L$.
+Figure 1: The number of divisions in the T-Re-An method of first example in our paper, where we take $\beta=30$ and $\Lambda=L$.
+For the quantum reweight-annealing scheme, we adopt a similar procedure as introduced above.
+For $Z(s_{k-1})/Z(s_{k})=\langle(s_{k-1}/s_{k})^{n_{k,\text{odd}}}\rangle_{s_{k}}$, we can expect that when $s\to 0$, the number of the operators of the bonds with odd indices goes to zero in Eq.
+(3).
+Therefore we can take $n_{k,\text{odd}}\approx\tilde{n}_{k,\text{odd}}=\Gamma_{k}\beta s_{k}L$ ($\Gamma_{k}\beta s_{k}L^{2}$) for 1D (2D) system, and we find this hypothetical linear relation on $s$ is convenient and has sufficiently good performance.
+The number of divisions in this case similarly scales with $\mathcal{O}(\Gamma\beta L)$ for some uniform $\Gamma_{k}=\Gamma$ if in 1D system.
+A.3 (III) Computing entanglement entropy
+One important application of our method is to calculate the entanglement entropy in conjunction with the Bisognano-Wichmann theorem, which offers an excellently approximated functional form to the entanglement Hamiltonian in systems exhibiting translational symmetry.
+We briefly introduce the functional form of 1D/2D system with periodic boundary condition here.
+Figure 2: Illustration of the LBW Hamiltonian with an example of a 2D square lattice with size $2L\times L$ under half bipartition.
+The $d$ in Eq.
+(14) denotes the distance from the corresponding site/bond to the boundary (the orange dashed line).
+The spacing of lattice is taken to be $1$.
+As it shown in Fig. 2, for an 1D/2D system under half bipartition (cornerless) with $A$ and $B$ part, if the Hamiltonian takes the form
+$\begin{split}H=&\Gamma\sum_{x,y,\delta=\pm 1}[h_{(x,y),(x+\delta,y)}+h_{(x,y),(x,y+\delta)}]\\ &+\Theta\sum_{x,y}l_{(x,y)}\\ \end{split}$ (12)
+with nearest-neighbor coupling $\Gamma$ and on-site coupling $\Theta$, then the corresponding lattice Bisognano-Wichmann (LBW) Hamiltonian $\tilde{H}_{A}$ is
+$\begin{split}\tilde{H}_{A}=&\frac{2\pi}{v}\bigg{[}\sum_{x,y,\delta=\pm 1}[\Gamma_{x}h_{(x,y),(x+\delta,y)}+\Gamma_{y}h_{(x,y),(x,y+\delta)}]\\ &+\Theta_{x}\sum_{x,y}l_{(x,y)}\bigg{]}\end{split}$ (13)
+where $v$ denotes the sound velocity and $\Gamma$, $\Theta$ are modified to be $\Gamma_{x},\ \Gamma_{y}$, and $\Theta_{x}$. For period boundary condition, we have
+$\begin{split}\Gamma_{\delta=x,y}&=\frac{d(L-d)}{L}\\ \Theta_{x}&=\frac{d(L-d)}{L}\\ \end{split}$ (14)
+which depend on the distance $d$ to the boundary of the corresponding bond/site term (See Fig. 2).
+With the LBW Hamiltonian $\tilde{H}_{A}$, we can define the LBW reduced density matrix $\tilde{\rho}_{A}$, which is
+$\tilde{\rho}_{A}\equiv\frac{e^{-\tilde{H}_{A}}}{\text{tr}\{e^{-\tilde{H}_{A}}\}}\approx\rho_{A}$ (15)
+where $\rho_{A}$ is the real reduced density matrix. Therefore we can compute the Von Neumann entropy $S(\rho_{A})$ by
+$S(\rho_{A})\approx-\text{tr}\{\tilde{\rho}_{A}\ln\tilde{\rho}_{A}\}=\langle\tilde{H}_{A}\rangle+\ln\tilde{Z}_{A}$ (16)
+where $Z_{A}\equiv\text{tr}\{e^{-\tilde{H}_{A}}\}$.
+Attention that Eq.
+(16) is exactly the thermal entropy of $\tilde{H}_{A}$ at $\beta=1$.
+Since it requires the value of $\ln\tilde{Z}_{A}$, our method can thus be used to calculate the Von Neumann entropy of lattice systems with translational symmetry.
+In addition to the LBW-HM model shown in the main body of the paper, we present another example of the 1D transverse field Ising model (TFIM), and calculate its Von Neumann entropy at the critical point with the reweight-annealing method.
+The exact result is $S(L)\propto(c/3)\ln L$, where $c=0.5$ is the central charge, and our method gives $c\approx 0.501(4)$, as it shown in Fig.
+3.
+Figure 3: For the 1D TFIM model at the critical point under PBC, the Von Neumann entropy obtained via calculating the thermal entropy of the corresponding BW-TFIM model at $\beta=1$.
+The central charge extracted from our Re-An simulations is $c=0.501(4)$, which is consistent with the theoretical result $0.5$.

### v5

@@
-Reweight-annealing method for calculating the value of partition function via quantum Monte Carlo
+Reweight-annealing method for evaluating the partition function via quantum Monte Carlo calculations
 Abstract
@@
 Here we present an unbiased but low-technical-barrier algorithm within the quantum Monte Carlo framework, which has exceptionally high accuracy and no systemic error.
-Compared with the conventional specific heat integral method and Wang-Landau sampling algorithm , our method can obtain a much more accurate result of the sub-leading coefficient of the entropy consistent with CFT.
+Compared with the conventional specific heat integral method and Wang-Landau sampling algorithm, our method can obtain a much more accurate result of the sub-leading coefficient of the entropy consistent with theory.
 This method can be widely used in both classical and quantum Monte Carlo simulations and is easy to be parallelized on computer.
+I Introduction
 […]
@@
 For phase transitions no matter within or beyond the Ginzburg-Landau paradigm, an efficient and accurate way to straightforwardly obtain PF would be crucial, because its (high order) derivative reveals potential phase transitions.
-The thermal entropy also relates to the distinctiveness of many novel condensed matters, such as spin ice Snyder et al.
-(2001); Skjærvø et al.
-(2020); Udagawa et al.
-(2021); Bramwell and Harris (2020); Udagawa et al.
-(2021), topological order Jiang et al.
-(2012); Castelnovo et al.
-(2012); Sachdev (2018); Levin and Wen (2006); Yan et al.
-(2021, 2022), fracton phases Pretko et al.
-(2020); Nandkishore and Hermele (2019); Gromov and Radzihovsky (2024); Zhou et al.
-(2022), etc.
+The thermal entropy also relates to the distinctiveness of many novel condensed matters, such as spin ice [20, 21, 22, 23, 22, 24], topological order [25, 26, 27, 28, 29, 30], fracton phases [31, 32, 33, 34], etc.
 Besides, the precision of entropy can even reveal important intrinsic physics.
-In conformal field theory (CFT), the prefactor in front of the sub-leading term of entropy often carries universal information Calabrese and Lefevre (2008); Fradkin and Moore (2006); Nussinov and Ortiz (2009a, b); Casini and Huerta (2007); Ji and Wen (2019, 2020); Kong et al.
-(2020); Wu et al.
-(2021a); Zhao et al.
-(2021); Wu et al.
-(2021b).
+In conformal field theory (CFT), the prefactor in front of the sub-leading term of entropy often carries universal information [35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45].
 Even a small numerical error may make our concerned quantities deviate a lot because the sub-leading term is much more sensitive than the leading one.
 In this paper, we present a more accurate and efficient method for calculating PF, dubbed reweight-annealing algorithm, within the framework of Monte Carlo (MC) simulation, either quantum or classical.
-It is inspired by thermal and quantum annealing algorithm in the quantum simulation area Huse and Fisher (1986); Santoro et al.
-(2002); Lucas (2014); Kadowaki and Nishimori (1998); Santoro et al.
-(2002); Mézard and Montanari (2009); Heim et al.
-(2015); Yan et al.
-(2023); Ding et al.
-(2023) and recent developments of the incremental trick in the entanglement calculations of QMC Alba (2017); D’Emidio (2020); Zhao et al.
-(2022a, b); Pan et al.
-(2023); Zhou et al.
-(2024).
-Our method not only has high accuracy for calculating PF, but is unbiased, easy to implement and naturally can be parallelized on computer.
-Basic idea.- Given a partition function $Z({\mathbf{p}})$, where ${\mathbf{p}}=[p_{1},p_{2},\cdots]$ denotes the parameters in the PF such as the temperature $T$ and some Hamiltonian parameter $s$.
+It is inspired by thermal and quantum annealing algorithm in the quantum simulation area [46, 47, 48, 49, 47, 50, 51, 52, 53, 54] and recent developments of the incremental trick in the entanglement calculations of QMC [55, 56, 57, 58, 59, 60], while the earliest idea can date back to the thermodynamic integration [61, 62, 63], and many of its variants found the applications in both condensed matter physics and high energy physics [64, 65, 66, 67, 68, 69, 70].
+The method not only has high accuracy for calculating PF, but is unbiased, easy to implement and naturally can be parallelized on computer.
+With the annealing scheme we introduce below, a polynomial complexity can be proved, enabling better control over the computational precision for large systems.
+This paper is arranged as follows.
+In Sec.
+II, we elaborate the algorithmic and technical details of our method, including the annealing scheme, then to demonstrate the efficacy of the method, we provide three examples in Sec.
+III.
+In the end, a summary and discussions are put in Sec.
+IV.
+II Method
+II.1 Basic idea
+Given a partition function $Z(\vec{p})$, where $\vec{p}=[p_{1},p_{2},\cdots]$ denotes the parameters in the PF such as the temperature $T$ and some Hamiltonian parameter $s$.
 For example, Fig.1 (a) shows a diagram of the two parameters.
-For two different parameter points ${\mathbf{p}}^{\prime}$ and ${\mathbf{p}}^{\prime\prime}$, we define the reweighting ratio which can be measured in Monte Carlo
-$\frac{Z({\mathbf{p}}^{\prime})}{Z({\mathbf{p}}^{\prime\prime})}=\bigg{\langle}\frac{W({\mathbf{p}}^{\prime})}{W({\mathbf{p}}^{\prime\prime})}\bigg{\rangle}_{{\mathbf{p}}^{\prime\prime}}$ (1)
-where $W({\mathbf{p}}^{\prime})$ and $W({\mathbf{p}}^{\prime\prime})$ denote the weight of the same configuration under different parameters ${\mathbf{p}}^{\prime}$ and ${\mathbf{p}}^{\prime\prime}$, and the subscript indicates the result comes from the MC simulation under parameter ${\mathbf{p}}^{\prime\prime}$.
-As an example, we take ${\mathbf{p}}=\beta$, where $\beta$ denotes the inverse temperature.
-Under the framework of stochastic series expansion (SSE), we can then derive $Z(\beta^{\prime})/Z(\beta^{\prime\prime})=\langle(\beta^{\prime}/\beta^{\prime\prime})^{n}\rangle_{\beta^{\prime\prime}}$, where $n$ is the number of non-identity operators in the SSE simulation.
-The details of the $Z(\beta^{\prime})/Z(\beta^{\prime\prime})$ have been explained in the supplemental materials (SM).
+For two different parameter points $\vec{p}^{\prime}$ and $\vec{p}^{\prime\prime}$, we define the reweighting ratio which can be measured in Monte Carlo
+$\frac{Z(\vec{p}^{\prime})}{Z(\vec{p}^{\prime\prime})}=\bigg{\langle}\frac{W(\vec{p}^{\prime})}{W(\vec{p}^{\prime\prime})}\bigg{\rangle}_{\vec{p}^{\prime\prime}}$ (1)
+where $W(\vec{p}^{\prime})$ and $W(\vec{p}^{\prime\prime})$ denote the weight of the same configuration under different parameters $\vec{p}^{\prime}$ and $\vec{p}^{\prime\prime}$, and the subscript indicates the result comes from the MC simulation under parameter $\vec{p}^{\prime\prime}$.
+As an example, we take $\vec{p}=\beta$, where $\beta$ denotes the inverse temperature.
+Under the framework of stochastic series expansion (SSE)[1, 2], we can then derive $Z(\beta^{\prime})/Z(\beta^{\prime\prime})=\langle(\beta^{\prime}/\beta^{\prime\prime})^{n}\rangle_{\beta^{\prime\prime}}$, where $n$ is the number of non-identity operators in the SSE simulation.
+The details of the $Z(\beta^{\prime})/Z(\beta^{\prime\prime})$ can be found in Appx.
+A.
 Figure 1: Schematic diagram of the reweight-annealing method.
@@
 If these two distributions are close to each other, the effect of the reweighting would be good since the importance sampling can approximately be kept.
-The second step is to identify a reference point ${\mathbf{p}}_{0}$, of which the value of $Z({\mathbf{p}}_{0})$ has already been known.
-Therefore, by evaluating $Z({\mathbf{p}}_{0})/Z({\mathbf{p}}^{\prime\prime})$, we can directly determine the value of $Z({\mathbf{p}}^{\prime\prime})$.
-In the example that ${\mathbf{p}}=\beta$, apparently $Z(0)=d^{N}$ is pre-knowledge because all the configurations are equally important at infinite high temperature, where $N$ and $d$ are the number of particles and the degree of freedom of a single particle respectively.
-However, when ${\mathbf{p}}_{0}$ and ${\mathbf{p}}^{\prime\prime}$ are far away in the parameter space, the value of $Z({\mathbf{p}}_{0})/Z({\mathbf{p}}^{\prime\prime})$ would be close to zero or infinity, resulting in the problem of inefficiency and low precision for the MC simulation.
+The second step is to identify a reference point $\vec{p}_{0}$, of which the value of $Z(\vec{p}_{0})$ has already been known.
+Therefore, by evaluating $Z(\vec{p}_{0})/Z(\vec{p}^{\prime\prime})$, we can directly determine the value of $Z(\vec{p}^{\prime\prime})$.
+In the example that $\vec{p}=\beta$, apparently $Z(0)=d^{N}$ is pre-knowledge because all the configurations are equally important at infinite high temperature, where $N$ and $d$ are the number of particles and the degree of freedom of a single particle respectively.
+However, when $\vec{p}_{0}$ and $\vec{p}^{\prime\prime}$ are far away in the parameter space, the value of $Z(\vec{p}_{0})/Z(\vec{p}^{\prime\prime})$ would be close to zero or infinity, resulting in the problem of inefficiency and low precision for the MC simulation.
 This can be easily understood in the reweighting frame as shown in Fig.
 1 (b).
-If we want to use a well-known distribution $Z({\mathbf{p}}^{\prime\prime})=\sum W({\mathbf{p}}^{\prime\prime})$ to calculate another distribution $Z({\mathbf{p}}^{\prime})=\sum W({\mathbf{p}}^{\prime})$ via resetting the weight of samplings, the weight before/after the resetting of the same configuration should be close to each other.
-It is still an importance sampling in this sense and requires that ${\mathbf{p}}^{\prime}$ and ${\mathbf{p}}^{\prime\prime}$ should be close enough.
+If we want to use a well-known distribution $Z(\vec{p}^{\prime\prime})=\sum W(\vec{p}^{\prime\prime})$ to calculate another distribution $Z(\vec{p}^{\prime})=\sum W(\vec{p}^{\prime})$ via resetting the weight of samplings, the weight before/after the resetting of the same configuration should be close to each other.
+It is still an importance sampling in this sense and requires that $\vec{p}^{\prime}$ and $\vec{p}^{\prime\prime}$ should be close enough [62, 71].
 Based on this consideration, practically, we divide the calculation of of Eq. (1) into many steps to enhance precision, i.e.
-$\frac{Z({\mathbf{p}}_{0})}{Z({\mathbf{p}}^{\prime\prime})}=\prod_{k=1}^{m}\frac{Z({\mathbf{p}}_{k-1})}{Z({\mathbf{p}}_{k})}$ (2)
-where ${\mathbf{p}}_{m}\equiv{\mathbf{p}}^{\prime\prime}$ and $Z({\mathbf{p}}_{k-1})/Z({\mathbf{p}_{k}})\to 1$ as long as ${\mathbf{p}}_{k}\to{\mathbf{p}}_{k-1}$.
-Here, ${\mathbf{p}}$ can be perceived to gradually change from the point of interest ${\mathbf{p}}^{\prime\prime}$ to the reference point ${\mathbf{p}}_{0}$ along a trajectory (annealing path) in the parameter space, akin to an adiabatic process or annealing as shown in Fig.1 (a).
+$\frac{Z(\vec{p}_{0})}{Z(\vec{p}^{\prime\prime})}=\prod_{k=1}^{m}\frac{Z(\vec{p}_{k-1})}{Z(\vec{p}_{k})}$ (2)
+where $\vec{p}_{m}\equiv\vec{p}^{\prime\prime}$ and $Z(\vec{p}_{k-1})/Z(\vec{p_{k}})\to 1$ as long as $\vec{p}_{k}\to\vec{p}_{k-1}$.
+Here, $\vec{p}$ can be perceived to gradually change from the point of interest $\vec{p}^{\prime\prime}$ to the reference point $\vec{p}_{0}$ along a trajectory (annealing path) in the parameter space, akin to an adiabatic process or annealing as shown in Fig.1 (a).
 Therefore by calculating Eq.
-(2), not only we are able to achieve a higher precision of $Z({\mathbf{p}}^{\prime\prime})$, but concurrently track the variation of PF along this annealing path.
+(2), not only we are able to achieve a higher precision of $Z(\vec{p}^{\prime\prime})$, but concurrently track the variation of PF along this annealing path.
 This allows us to further analyze the variation of the free energy (and its derivatives) of the system.
-Apparently, the annealing path that ${\mathbf{p}}$ follows determines how large $m$ should we need and how ${\mathbf{p}}$ varies to perform the calculations of Eq.
-(2).
-It can be expected that, similar to conventional thermal annealing or quantum annealing, should the annealing trajectory approach a critical point, an increased number of divisions would be necessary in its vicinity.
-However, we point out that this actually affects little to our algorithm since numerical simulations only deal with finite system, where the criticality does not result in significant inefficiency.
-We have also verified this point in our simulations.
-For example, in the 2D spin-$1/2$ XXZ model which we present below, where ${\mathbf{p}}=T$ is the temperature, we implemented an equal division strategy with step size $\Delta T=0.01$.
-With $20000$ Monte Carlo samples for each bin, we are able to achieve a sufficiently smooth curve for further numerical differentiation.
-More discussions about the segmentation are in the section of “Dynamical annealing process”.
-Thermal and quantum reweight-annealing- As it shown in Fig.
-1, if ${\mathbf{p}}=\beta$, as previously mentioned, the annealing path connects an unknown inverse temperature point $\beta^{\prime\prime}$ to the infinite temperature point $\beta_{0}$, which serves as the reference.
+II.2 Thermal and quantum reweight-annealing
+As it shown in Fig.
+1, if $\vec{p}=\beta$, as previously mentioned, the annealing path connects an unknown inverse temperature point $\beta^{\prime\prime}$ to the infinite temperature point $\beta_{0}$, which serves as the reference.
 For convenience, we call this kind of process the thermal reweight-annealing (T-Re-An).
-Similarly, we can consider the so-called quantum reweight-annealing (Q-Re-An) by setting ${\mathbf{p}}=s$, where $s$ comes from $H(s)=sH_{0}+(1-s)H_{1}$, which is the archetypal Hamiltonian of quantum annealing with $Z(s)=e^{-\beta H(s)}$.
+Similarly, we can consider the so-called quantum reweight-annealing (Q-Re-An) by setting $\vec{p}=s$, where $s$ comes from $H(s)=sH_{0}+(1-s)H_{1}$, which is the archetypal Hamiltonian of quantum annealing with $Z(s)=e^{-\beta H(s)}$.
 Here, $H_{1}$ correspond to our reference point $s=0$, and we would like to obtain the PF related to $H_{0}$.
-We emphasis that different from the thermal version, the choice of $H_{1}$ here is much more flexible.
-For example, we can take $H_{1}=\sum_{i}\sigma_{i}^{x}$.
+Different from the thermal version, the choice of $H_{1}$ here is much more flexible.
+For instance, we can take $H_{1}=\sum_{i}\sigma_{i}^{x}$.
 Then at the zero temperature ($\beta\to\infty$), $H_{1}$ only has a unique ground state, thus its thermal entropy is minimized to be zero, resulting in $\ln Z(0)=-\lim_{\beta\to\infty}\beta E$, where $E=-N$ is the ground state energy of $H_{1}$.
 More details about the T-Re-An and Q-Re-An scheme can be found in the SM.
-Dynamical annealing process.- For the thermal and quantum reweight-annealing schemes, we provide a dynamical strategy to automatically regulate the alteration of ${\mathbf{p}}$ throughout the annealing process.
+II.3 Automatic annealing process
+Here we provide a scheme to automatically regulate the alteration of $\vec{p}$ throughout the annealing process.
 This not only serves as an economic way for the division in Eq.
-(2), but can avoid the problem of precision loss when different $r({\mathbf{p}}_{k-1},{\mathbf{p}}_{k})$ differ by several orders of magnitude in the computer.
-For demonstration, we only discuss the thermal case, i.e.
-${\mathbf{p}}=\beta$, in this section.
-Without loss of generality, we assume $\beta_{k-1}<\beta_{k}$ in Eq. (2), and define the rate $\alpha_{k}\equiv\beta_{k-1}/\beta_{k}<1$. Within the framework of SSE, we have
-$\frac{Z(\beta_{k-1})}{Z(\beta_{k})}=\langle\alpha^{n_{k}}_{k}\rangle_{\beta_{k}}$ (3)
-where $n_{k}$ is the number of all non-identity operators in the SSE simulation, and the subscript $\beta_{k}$ denotes that this simulation is performed at the parameter point $\beta_{k}$.
-The annealing process starts from $k=m$, and we require $Z(\beta_{k-1})/Z(\beta_{k})=\epsilon<1$, where the choice of $\epsilon$ should match $N_{\text{MC}}$, the number of MC steps we use for samplings, as the error in MC simulations is proportional to $1/\sqrt{N_{\text{MC}}}$.
-Next, we set $n_{k}\approx\tilde{n}_{k}=\Lambda_{k}\beta_{k}L$ ($\Lambda_{k}\beta_{k}L^{2}$) for 1D (2D) system, and $\Lambda_{k}$ can be ascertained through the examination of result convergence.
-This comes from the fact that in SSE, the energy satisfies $E_{k}=-\langle n_{k}\rangle/\beta_{k}$.
-Practically, we set $\Lambda_{k}$ to be the same for all $k$ for convenience.
-This strategy not only makes $Z(\beta_{k-1})/Z(\beta_{k})\approx\epsilon$ for all $k$, but determines the next step in the annealing path, i.e.
-$\beta_{k-1}=\alpha_{k}\beta_{k}$.
-The division will be more sparse when $\beta$ gets smaller as we expected, where the thermal fluctuation has destroyed any order and $Z(\beta_{k-1})/Z(\beta_{k})\to 1$, which has little contribution to the result.
-Apparently, this strategy naturally allows for parallelization of evaluating different $Z(\beta_{k-1})/Z(\beta_{k})$ in Eq.
+(2), but can avoid the problem of precision loss when different $Z(\vec{p}_{k-1})/Z(\vec{p}_{k})$ differ by several orders of magnitude in the computer.
+For demonstration, we discuss the thermal case, i.e.
+$\vec{p}=\beta$ in this section, and the case that $\vec{p}=s$ is similar.
+For $\beta_{0}<\beta$, when $\beta_{0}$ and $\beta$ are far away in the parameter space, to ensure importance sampling, we divide the interval $[\beta_{0},\beta]$ into $m$ subintervals $[\beta_{0},\beta_{1}]$, $[\beta_{1},\beta_{2}]$, $\cdots$, $[\beta_{m-1},\beta_{m}]$, where $\beta_{m}\equiv\beta$, i.e.
+$\frac{Z(\beta_{0})}{Z(\beta)}=\prod_{k=1}^{m}\frac{Z(\beta_{k-1})}{Z(\beta_{k})}$ (3)
+as Eq.
 (2).
-It is also easy to calculate that the number of divisions is around $(\beta-\beta_{0})\Lambda L/|\ln\epsilon|$ for large $L$ or $(\beta-\beta_{0})$, where $\Lambda_{k}=\Lambda$ for all $k$.
-Therefore, this dynamical annealing strategy does not require exponentially increasing number of divisions.
+Each subinterval $[\beta_{k-1},\beta_{k}]$ ought to be sufficiently narrow so that the distribution at $\beta_{k}$ can well samples the configurations at $\beta_{k-1}$.
+Otherwise, the sampling process would be inefficient.
+Then the natural questions following are how to partition $[\beta_{0},\beta]$ and the how many subintervals we need for Eq.
+(3).
+One crude way is just dividing $[\beta_{0},\beta]$ into equal subintervals or in a geometric progression.
+As long as the division number is large enough, the result would be accurate.
+However, this way is not economic.
+We can expect that when $\beta\to 0$ (approaching infinite high temperature), the division can be more sparse compared to that at low temperatures.
+Another worry is that if $\beta_{k}$ and $\beta_{k-1}$ are separated on the two sides of the critical point, we may need a bulk of samples to estimate $Z(\beta_{k-1})/Z(\beta_{k})$.
+In addition to the division problem, $Z(\beta_{0})/Z(\beta)$ could be an exponentially small number exceeding the data range of computer (e.g.
+$1.7\times 10^{-308}\sim 1.7\times 10^{308}$ for double type data), and we should store $\ln[Z(\beta_{0})/Z(\beta)]$ in the computer.
+However, simply applying the logarithm to Eq.
+(3) and converting the multiplication to addition may lead to precision loss in computer calculations if $\ln[Z(\beta_{k-1})/Z(\beta_{k})]$ and $\ln[Z(\beta_{k})/Z(\beta_{k+1})]$ differ by several orders of magnitude.
+To address the two problems above, we present the automatic/pseudo-automatic annealing scheme.
+The automatic annealing scheme starts by fixing $Z(\beta_{k-1})/Z(\beta_{k})=\epsilon<1$ for all $k$.
+We will talk about the choice of $\epsilon$ later, and first assume we have already selected some $\epsilon$.
+Recall that in SSE, the energy estimator is $E(\beta_{k})=-{\langle n\rangle_{\beta_{k}}}/{\beta_{k}}$, where $n_{k}\equiv\langle n\rangle_{\beta_{k}}$ denotes the average number of non-identity operators at temperature $\beta_{k}$, which itself is also an observable that can be estimated in SSE simulations.
+After we have achieved an estimation $\tilde{n}_{k}$ of $n_{k}$ from simulations (this is exactly same with measuring energy by counting the number of non-identity operators in each SSE configuration), we can bring the result into $Z(\beta_{k-1})/Z(\beta_{k})$ to have
+$\epsilon=\frac{Z(\beta_{k-1})}{Z(\beta_{k})}=\bigg{\langle}\bigg{(}\frac{\beta_{k-1}}{\beta_{k}}\bigg{)}^{n}\bigg{\rangle}_{\beta_{k}}\approx\bigg{(}\frac{\beta_{k-1}}{\beta_{k}}\bigg{)}^{\tilde{n}_{k}}$ (4)
+Attention that Eq. (4) is just an approximation, but it gives us a way to divide $[\beta_{0},\beta]$.
+Imagine that we start from the $\beta\equiv\beta_{m}$, which is the right end point of $[\beta_{0},\beta]$, then using Eq. (4), we can determine the position (value) of $\beta_{m-1}$, which is
+$\beta_{m-1}=\epsilon^{1/\tilde{n}_{k}}\beta_{m}$ (5)
+Similarly, we can obtain $\beta_{m-2},\beta_{m-3},\cdots$, and it stops when we encounter some $\beta_{j}$ such that $\beta_{j-1}<\beta_{0}$.
+Then $\beta_{j}$ is set to be $\beta_{1}$ and the division procedure is over.
+Notice that each $\beta_{k-1}$ must rely on the knowledge of $\beta_{k}$, and each $\tilde{n}_{k}$ is obtained from the corresponding SSE simulations.
+This division procedure is similar to a standard (thermal or quantum) annealing process, in which we start from an initial parameter point, and step-by-step change it to a target parameter point.
+Here we go from $\beta\equiv\beta_{m}$ to $\beta_{0}$.
+Of course one can inversely go from $\beta_{0}$ to $\beta$ by considering $Z(\beta)/Z(\beta_{0})$ rather than $Z(\beta_{0})/Z(\beta)$.
+About the choice of $\epsilon$, since we have to estimate $Z(\beta_{k-1})/Z(\beta_{k})$, we hope this value smaller than $1$ but close to $1$ so that the distribution at $\beta_{k-1}$ and $\beta_{k}$ would have big overlap to make the simulations more efficient.
+Just as we perform some common MC simulations to estimate some observable, the MC steps and the number of bins required for an accurate estimation should be tried.
+Here in our algorithm, it is similar actually.
+After we fix some $\epsilon$, we have to determine the required MC steps and bins via some trial simulations.
+We can expect that if we take a very small $\epsilon$, the MC steps required for an accurate estimation would be large.
+However, no matter what $\epsilon$ we choose, in principle, as long as we have enough MC steps and bins, the estimation of $Z(\beta_{k-1})/Z(\beta_{k})$ can be accurate.
+If $\epsilon$ is very close to $1$, we would also have some problems because the number of divided segments $m$ would be very large.
+Therefore there is some tradeoff about the choice of $\epsilon$.
+II.4 Pseudo-automatic annealing process
+The “automatic annealing” scheme mentioned above can not be parallelized since we have to achieve each $\tilde{n}_{k}$ every time from the simulations so that we can go to the next point.
+Therefore, practically, we give a priori estimate of the value of $\tilde{n}_{k}$ without doing the simulations.
+This is what we call the pseudo-automatic annealing scheme used in our paper.
+Recall that energy is an extensive quantity, then it should satisfy $E\sim L^{d}$ for a $d$-dimensional system.
+In SSE, we then have $\langle n\rangle_{\beta_{k}}\sim\beta_{k}L^{d}$.
+Here $\beta_{k}$, $L$ and $d$ are given, therefore to obtain $\langle n\rangle_{\beta_{k}}$, we need to know some prefactor $\Lambda_{k}$ such that $\langle n\rangle_{\beta_{k}}\simeq\Lambda_{k}\beta_{k}L^{d}\approx\tilde{n}_{k}$.
+It might be questioned that not even we have to decide the value of $\epsilon$, but we have another hyperparameter $\Lambda_{k}$ to determine.
+In fact, we still have one since now Eq.
+(4) becomes
+$\epsilon\approx\bigg{(}\frac{\beta_{k-1}}{\beta_{k}}\bigg{)}^{\Lambda_{k}\beta_{k}L^{d}}$ (6)
+If we substitute the real $\Lambda_{k}$ with some incorrect $\tilde{\Lambda}_{k}$ we manually set, this just effectively changes the value of $Z(\beta_{k-1})/Z(\beta_{k})$ from $\epsilon$ to some $\epsilon_{\text{eff}}$.
+Therefore we can fix $\Lambda_{k}$ to be some value such as $L$ in our simulations for all $k$ for convenience, then the problem reduces to choosing a good $\epsilon$.
+The uniform choice of $\Lambda_{k}\equiv\Lambda$ would make different $Z(\beta_{k-1})/Z(\beta_{k})$ slightly deviated from our desired ratio $\epsilon_{\text{eff}}$, but it is okay since we always have to determine the MC steps and bins for estimation.
+In this aspect, there is no different no matter for $\epsilon_{\text{eff}}$ (with small deviations at each point) or the $\epsilon$ in the “automatic annealing” scheme in the last subsection.
+Next we show that this pseudo-automatic annealing scheme actually requires the number of divisions scales proportional to the system size and the inverse temperature for large systems.
+For the 1D system as an example,
+$\beta_{k-1}=\epsilon^{\frac{1}{\Lambda\beta_{k}L}}\beta_{k}\approx\bigg{(}1+\frac{\ln\epsilon}{\Lambda\beta_{k}L}\bigg{)}\beta_{k}$ (7)
+or
+$\beta_{k}-\beta_{k-1}=-\frac{\ln\epsilon}{\Lambda L}$ (8)
+then the number of divisions is around $(\beta-\beta_{0})\Lambda L/|\ln\epsilon|$.
+Notice that we have ignored the case that $\beta_{k}\to 0$ in Eq.(7).
+This is because when $\beta\to 0$, we have $\epsilon^{1/(\Lambda\beta_{k}L)}\to 0$ ($\epsilon<1$), then the division becomes sparser and has little contributions to the total number of divisions.
+Another thing worthy of attention is that when $\beta\to 0$, the small value of $\epsilon^{1/(\Lambda\beta_{k}L)}\to 0$ would also exceed the range of a number saved in the computer.
+Therefore we manually set a lower bound for $\alpha_{k}$ (e.g.
+$\min{\alpha_{k}}=\epsilon$ in our simulation).
+The existence of this lower bound makes the division in the region that $\beta\to 0$) slightly denser than that Eq.
+(6) requires.
+Since most $Z(\beta_{k-1})/Z(\beta_{k})$ still at the same order of magnitude, and the more divisions when $\beta\to 0$ also only constitutes a very small proportion.
+This affects little to the precision as well as the number of divisions calculated in Eq.
+(8).
+In Fig.
+2, we plot the number of divisions in the T-Re-An method of first example (see Sec.
+III.1) in our paper for reference, where we take $\Lambda=L$, and compare it with the estimation number $(\beta-\beta_{0})\Lambda L/|\ln(\epsilon)|$.
+Figure 2: The number of divisions in the T-Re-An method of first example (see Sec. III.1), where we take $\epsilon=0.01$, $\beta=30$ and $\Lambda=L$.
+III Examples
+III.1 1D spin-1/2 Heisenberg chain
+As the first example, we consider an 1D spin-1/2 antiferromagnetic (AFM) Heisenberg chain with periodic boundary condition (PBC) with $\beta=30$ and length $L$.
+We use both the T-Re-An method and the Q-Re-An method to calculate the PF.
+For the quantum version, we write
+$H(s)=s\bigg{[}\sum_{i}\mathbf{S}_{i}\cdot\mathbf{S}_{i+1}\bigg{]}+(1-s)\bigg{[}\sum_{i\in\text{even}}\mathbf{S}_{i}\cdot\mathbf{S}_{i+1}\bigg{]}$ (9)
+where $H_{0}=\sum_{i}\mathbf{S}_{i}\cdot\mathbf{S}_{i+1}$ and $H_{1}=\sum_{i\in\text{even}}\vec{S}_{i}\cdot\vec{S}_{i+1}$.
+Apparently, the unique ground state of $H_{1}$ is a tensor product of $L/2$ singlets, each carrying energy $-3/4$ ($\hbar=1$).
+The estimator for $Z(s^{\prime})/Z(s^{\prime\prime})$ is $\langle(s^{\prime}/s^{\prime\prime})^{n_{\text{odd}}}\rangle_{s^{\prime\prime}}$ in this model (see the derivation in Appx.
+A), with $n_{\text{odd}}$ denoting the amount of non-identity operators acting on all the bonds with odd indices within the framework of SSE.
 Figure 3: For $\beta=30$, the values of $\ln Z$ as a function of the chain length $L$, where both T-Re-An method and Q-Re-An method have a perfect match with ED.
@@
 Attention that here in practical SSE simulations, the Hamiltonian has a energy shift of $-1/4$ for each two-body term.
-1D spin-1/2 Heisenberg chain.- As the first example, we consider an 1D spin-1/2 antiferromagnetic (AFM) Heisenberg chain with periodic boundary condition (PBC) with $\beta=30$ and length $L$.
-We use both the T-Re-An method and the Q-Re-An method to calculate the PF.
-For the quantum version, we write
-$H(s)=s\bigg{[}\sum_{i}\mathbf{S}_{i}\cdot\mathbf{S}_{i+1}\bigg{]}+(1-s)\bigg{[}\sum_{i\in\text{even}}\mathbf{S}_{i}\cdot\mathbf{S}_{i+1}\bigg{]}$ (4)
-where $H_{0}=\sum_{i}\mathbf{S}_{i}\cdot\mathbf{S}_{i+1}$ and $H_{1}=\sum_{i\in\text{even}}{\mathbf{S}}_{i}\cdot{\mathbf{S}}_{i+1}$.
-Apparently, the unique ground state of $H_{1}$ is a tensor product of $L/2$ singlets, each carrying energy $-3/4$ ($\hbar=1$).
-The estimator for $Z(s^{\prime})/Z(s^{\prime\prime})$ is $\langle(s^{\prime}/s^{\prime\prime})^{n_{\text{odd}}}\rangle_{s^{\prime\prime}}$ in this model, with $n_{\text{odd}}$ denoting the amount of non-identity operators acting on all the bonds with odd indices within the framework of SSE.
+For the pseudo-annealing scheme, we adopt a similar procedure as in Sec.
+II.4.
+For $Z(s_{k-1})/Z(s_{k})=\langle(s_{k-1}/s_{k})^{n_{k,\text{odd}}}\rangle_{s_{k}}$, we can expect that when $s\to 0$, the number of the operators of the bonds with odd indices goes to zero.
+Therefore we can take $n_{k,\text{odd}}\approx\tilde{n}_{k,\text{odd}}=\Gamma_{k}\beta s_{k}L$ ($\Gamma_{k}\beta s_{k}L^{2}$) for 1D (2D) system, and we find this hypothetical linear relation on $s$ is convenient and has sufficiently good performance just as the first example shows.
+The number of divisions in this case similarly scales with $\mathcal{O}(\Gamma\beta L)$ for some uniform $\Gamma_{k}=\Gamma$ if in 1D systems.
 As it is shown in Fig.
-2, both the results of T-Re-An and Q-Re-An method can match that of the exact diagonalization (ED) method exceptionally well for small lengths, which verify the validity of our methods.
+3, both the results of T-Re-An and Q-Re-An method can match that of the exact diagonalization (ED) method exceptionally well for small lengths, which verify the validity of our methods.
 The results are also consistent with the expectation that $\ln Z\propto L+o(L)$, where $o(L)$ means higher order terms which is negligible compared with $L$.
@@
 Figure 4: (a) Thermodynamic entropies $S(L)$ as a function of $L$ calculated from different algorithms; (b) The $\ln L$ term extracted by calculating $2S(L)-S(2L)$, and the fitting slope (dashed green line) here is 1.00(2).
-2D LBW-HM.- To further exploit the advantage of our method, we calculate the thermodynamic entropy of a 2D lattice Bisognano-Wichmann Heisenberg model (LBW-HM) Giudici et al.
-(2018); Dalmonte et al.
-(2018), the thermodynamic entropy $S$ of the which at $\beta=1$ provides a good approximation for the Von Neumann entanglement entropy of a 2D PBC square lattice spin-1/2 AFM Heisenberg model under smooth bipartition Mendes-Santos et al.
-(2019, 2020).
-According to the CFT prediction and previous numerical results Metlitski and Grover (2015); Laflorencie (2016); Zhao et al.
-(2022a); Deng et al.
-(2023), the coefficient of sub-leading term of the entropy reflects the number of Goldstone modes.
-Here, the thermal entropy of the LBW-HM at $\beta=1$ should obey $S(L)=aL+b\ln L+c$, and $b=n_{G}/2$, where $n_{G}=2$ is the number of Goldstone modes of the 2D spin-1/2 AFM Heisenberg model Metlitski and Grover (2011); Laflorencie (2016).
+III.2 2D LBW-HM
+To further exploit the advantage of our method, we calculate the thermodynamic entropy of a 2D lattice Bisognano-Wichmann Heisenberg model (LBW-HM) [72, 73], the thermodynamic entropy $S$ of the which at $\beta=1$ provides a good approximation for the Von Neumann entanglement entropy of a 2D PBC square lattice spin-1/2 AFM Heisenberg model under smooth bipartition [74, 75].
+A brief review of this theory can be found in Appx.
+B.
+According to the CFT prediction and previous numerical results [76, 77, 57, 78], the coefficient of sub-leading term of the entropy reflects the number of Goldstone modes.
+Here, the thermal entropy of the LBW-HM at $\beta=1$ should obey $S(L)=aL+b\ln L+c$, and $b=n_{G}/2$, where $n_{G}=2$ is the number of Goldstone modes of the 2D spin-1/2 AFM Heisenberg model [79, 77].
 In this context, coarse result of entropy is disfavored because the the coefficient $b$ in front of the sub-leading term is sensitive to errors.
@@
 As it shown in Fig.
-3, the $C$-int method has great systemic error because the number of division is adequate for our method, but not the $C$-int method, indicating more computational resources needed for it to achieve better accuracy.
-For the WL algorithm, on another hand, the allocation of total computational resources should not be predetermined as it hinges on when we can achieve a converged and flat histogram.
-Fig.
-3 shows great error of the results obtained from the vanilla WL samplings Troyer et al.
-(2003).
-This reflects the problem of error saturation, which is intrinsic in this algorithm due to an approximately flat histogram and the empirical refinement parameter Belardinelli et al.
-(2008); Belardinelli and Pereyra (2007); Zhou and Bhatt (2005).
+4, the $C$-int method has great systemic error because the number of division is adequate for our method, but not the $C$-int method, indicating more computational resources needed for it to achieve better accuracy.
+For the WL algorithm, the number of density of states to be flattened in each histogram should be greater than $\beta E(\beta)$, where $E(\beta)$ is the energy at $\beta$, which is also a (roughly) polynomial complexity with the system size [15].
+However, the main problem of this algorithm is the so-called error saturation, which is intrinsic, preventing us from obtaining results of an arbitrary precision generally and making this algorithm basically biased [80, 81, 18, 19].
+In Fig.
+4, we show great error of the results obtained from the vanilla WL samplings [15], and to achieve an arbitrary precision within the Wang-Landau framework is not generally easy.
 Another disadvantage of the WL sampling compared with our method is that it cannot be parallelized: the two subroutines for flattening the corresponding histograms before and after an adjustment of the refinement parameter, must be executed sequentially.
-To achieve an arbitrary precision within the Wang-Landau framework is not easy.
-In contrast, our method is much more accessible as we only need to measure the observable of $Z({\mathbf{p}}_{k-1})/Z({\mathbf{p}}_{k})$ in many conventional and mature MC algorithms.
-2D XXZ model.- As the last example, we consider a much more complex model, a 2D spin-1/2 XXZ model with next nearest neighbor interaction,
+In contrast, our unbiased method is able to achieve an arbitrary precision, easy to parallelize, and much more accessible as we only need to measure the observable of $Z(\vec{p}_{k-1})/Z(\vec{p}_{k})$ in many conventional and mature MC algorithms.
+III.3 2D XXZ model
+As the last example, we consider a much more complex model, a 2D spin-1/2 XXZ model with next nearest neighbor interaction,
 $H=-\sum_{\langle i,j\rangle}[2S^{z}_{i}S^{z}_{j}+S^{x}_{i}S^{x}_{j}+S^{y}_{i}S^{y}_{j}]$
@@
 The peak diverges more when the size becomes larger, which probes the phase transition point here; (b) The Binder ratios in this model at different sizes intersecting at a point which is highly consistent with the peak obtained by Re-An algorithm.
-For convenience, here we use equal distance of $T=1/\beta=0.01$ to implement our Re-An method, and it is also easy for numerical differentiation.
-Because a constant factor of the PF does not affect the (higher order) derivative of $F$, thus we choose $Z(T^{*}=1.95)$ as the reference to replace annealing to a known point.
-The PF we gained is actually $Z(T)/Z(T^{*})$ where $Z(T^{*})$ is an unknown constant.
+Here we use equal division of $T=1/\beta=0.01$, which is convenient for further numerical differentiation.
+We found that with step size $\Delta T=0.01$ and $20000$ MC samples for each bin, a sufficiently smooth curve can be obtained.
+Because a constant factor of the PF does not affect the (higher order) derivative of $F$, we choose $Z(T^{*}=1.95)$ as the reference point rather than some known point.
+The PF we calculated is thus $Z(T)/Z(T^{*})$ where $Z(T^{*})$ is an unknown constant.
 Then we indeed observe a divergent peak of the second order derivative of the free energy as shown in Fig.
-4(a), which accurately probes the phase transition point, matching with the intersection point of Binder ratios $R_{2}=\langle M^{4}\rangle/\langle M^{2}\rangle^{2}$ of different sizes shown in Fig.
-4(b), where $M$ denotes the magnetization.
+5(a), which accurately probes the phase transition point, matching with the intersection point of Binder ratios $R_{2}=\langle M^{4}\rangle/\langle M^{2}\rangle^{2}$ of different sizes shown in Fig.
+5(b), where $M$ denotes the magnetization.
 This example also inspires us that connecting to a known point is not necessary when we just need the second or higher order derivative of free energy.
 The data also demonstrate the high-precision of our method, where the second order derivative is still smooth and precise.
-Summary.- We propose an unbiased algorithm within Monte Carlo framework, which has low-technical barrier and no systemic error.
+IV Summary and discussions
+We propose an unbiased algorithm within the Monte Carlo framework, which has low-technical barrier and no systemic error.
 The method can be easily implemented for parallel computation and can obtain the value of partition function, free energy and thermal entropy with high precision.
-The method has two keynotes: (1) Find a parameter path in whole space (either temperature or external parameter) connecting a known point and the to-be-solved one.
+The method has two keynotes: (1) Finding a parameter path in whole space (either temperature or external parameter) connecting a known point and the to-be-solved one.
 If we only want to calculate the (high order) derivative of the PF, even fixing an unknown point as the reference is enough; (2) Annealing along the path to calculate the final result through a propagation formula.
@@
 In a 2D XXZ model, the numerical second derivative of the Re-An data is still smooth and precise, which accurately probes the phase transition point.
-Our method make the efficient and high-precision calculation of entropy and other related physical quantities accessible, which can reveal some universal information, such as CFT, criticality and symmetry breaking.
-Acknowledgement.- We acknowledge the start-up funding of Westlake University.
+Our method makes the efficient and high-precision calculation of entropy and other related physical quantities accessible, which can reveal some universal information, such as CFT, criticality and symmetry breaking.
+The final thing worth of emphasizing is that the reweight-annealing method discussed here is a general paradigm, not limited in computing the standard partition function, but can be extended to any generalized partition function as long as the two keynotes can be satisfied.
+One representative example has been shown in [82], where the authors developed an algorithm to compute the Rényi entropy using this reweight-annealing framework.
+Acknowledgement
+We acknowledge the start-up funding of Westlake University.
 The authors also acknowledge Beijing PARATERA Tech Co.,Ltd.(https://www.paratera.com/) for providing HPC resources that have contributed to the research results reported within this paper.
-CC and ZY thank the support of the National Natural Science Foundation of China (grant nos.
-12174167, 12247101).
-Appendix A Supplemental material
-A.1 (I) Estimators in thermal and quantum reweight-annealing under the SSE framework
-For the thermal case, the estimator of $Z(\beta^{\prime})/Z(\beta^{\prime\prime})$ is general, which is
-$\begin{split}\frac{Z(\beta^{\prime})}{Z(\beta^{\prime\prime})}=&\frac{1}{Z(\beta^{\prime\prime})}\text{tr}\{e^{-\beta^{\prime}H}\}\\ =&\frac{1}{Z(\beta^{\prime\prime})}\text{tr}\bigg{\{}\sum_{n}\bigg{(}\frac{\beta^{\prime}}{\beta^{\prime\prime}}\bigg{)}^{n}\frac{(\beta^{\prime\prime})^{n}}{n!}(-H)^{n}\bigg{\}}\\ =&\bigg{\langle}\bigg{(}\frac{\beta^{\prime}}{\beta^{\prime\prime}}\bigg{)}^{n}\bigg{\rangle}_{\beta^{\prime\prime}}\\ \end{split}$ (1)
-and the subscript denotes that this simulation should be performed at the inverse temperature $\beta^{\prime\prime}$.
-For the quantum case with $H(s)=sH_{0}+(1-s)H_{1}=H_{1}+s(H_{0}-H_{1})$, $Z(s^{\prime})/Z(s^{\prime\prime})$ usually takes the form $\langle(s^{\prime}/s^{\prime\prime})^{n_{s}}\rangle_{s^{\prime\prime}}$ within the framework of SSE, where $n_{s}$ is the number of non-identity operators related to $(H_{0}-H_{1})$.
+CC and ZY thank the support of the National Natural Science Foundation of China (grant no.
+12247101).
+NM thanks the support of the National Natural Science Foundation of China (grant no.
+12004020).
+GP thanks the Würzburg-Dresden Cluster of Excellence on Complexity and Topology in Quantum Matter - ct.qmat (EXC 2147, Project No.
+390858490).
+Note: After we finished this work, we found a related work posted on arXiv two weeks ago [83].
+Appendix A Estimators in thermal and quantum reweight-annealing under the SSE framework
+We first review the reweighting trick in a general MC method with the inverse temperature $\beta$ for illustrations (the thermal case), then we will return to the SSE method, which is just one way of transforming the quantum degrees of freedom to classical ones in the partition function.
+Consider a partition function $Z(\beta)=\sum_{C}W(\beta;C)$, where the weights $\{W(\beta;C)\}$ are non-negative real numbers (such that we can perform MC simulations) and $C$ denotes each configuration.
+For the expectation value of some diagonal operator $O$ at the inverse temperature $\beta^{\prime\prime}$, This can be achieved as follows:
+$\begin{split}\langle O\rangle_{\beta^{\prime\prime}}=&\frac{\sum_{C}W(\beta^{\prime\prime};C)O(C)}{Z(\beta^{\prime\prime})}\\ \end{split}$ (11)
+where $O(C)$ is the value of the operator $O$ under the configuration $C$. In our paper, a specific form of the operator $O=[W(\beta^{\prime})/W(\beta^{\prime\prime})]$ is discussed, then we have
+$\begin{split}\bigg{\langle}\frac{W(\beta^{\prime})}{W(\beta^{\prime\prime})}\bigg{\rangle}_{\beta^{\prime\prime}}=&\frac{1}{Z(\beta^{\prime\prime})}\sum_{C}W(\beta^{\prime\prime};C)\frac{W(\beta^{\prime};C)}{W(\beta^{\prime\prime};C)}\\ =&\frac{\sum_{C}W(\beta^{\prime};C)}{Z(\beta^{\prime\prime})}\\ =&\frac{Z(\beta^{\prime})}{Z(\beta^{\prime\prime})}\end{split}$ (12)
+Eq.
+(12) is the key relation in this paper, in which we sample the configurations with probabilities $p(\beta^{\prime\prime};C)=W(\beta^{\prime\prime};C)/Z(\beta^{\prime\prime})$ at inverse temperature $\beta^{\prime\prime}$, but recalculating the weights of the sampled configuration $C$ at another inverse temperature $\beta^{\prime}$, i.e., the $W(\beta^{\prime};C)$.
+What Eq.
+(12) indicates is that through averaging the weight ratio of each configuration, the ratio of partition functions at different inverse temperature $Z(\beta^{\prime})/Z(\beta^{\prime\prime})$ can be obtained.
+Now we turn to SSE, the demonstrative QMC method used in this paper [1, 2]. Apply the Taylor expansion, we have
+$Z(\beta)=\text{tr}\bigg{\{}\sum_{n}\frac{\beta^{n}}{n!}(-H)^{n}\bigg{\}}=\sum_{n}\frac{\beta^{n}}{n!}\text{tr}\{(-H)^{n}\}$ (13)
+Then we write $H=-\sum_{a,b}H_{a,b}$, where $H_{a,b}$ are some non-identity local operator with $a$ denoting the type of it (diagonal or off-diagonal under the simulating basis) and $b$ denotes the index of the space degree (bond or site) that the local operator is acting on.
+Hence
+$(-H)^{n}=\sum_{S_{n}}\prod_{p=1}^{n}H_{a_{p},b_{p}}$ (14)
+where $S_{n}$ denotes that we sum over all possible operator strings, and each of them is a product of $n$ local operators. Then we have
+$Z(\beta)=\sum_{n}\sum_{S_{n}}\frac{\beta^{n}}{n!}\text{tr}\bigg{\{}\prod_{p=1}^{n}H_{a_{p},b_{p}}\bigg{\}}$ (15)
+The important thing here is that because the product of $n$ local operators are interpreted as an operator string, therefore each string $S_{n}$ has exactly $n$ non-identity operators.
+In other words, each $S_{n}$ has the length $n$.
+Replacing the trace operation with a set of complete basis $\{\alpha_{0}\}$, we have
+$Z(\beta)=\sum_{n}\sum_{S_{n}}\sum_{\alpha_{0}}\frac{\beta^{n}}{n!}\bigg{\langle}\alpha_{0}\bigg{|}\prod_{p=1}^{n}H_{a_{p},b_{p}}\bigg{|}\alpha_{0}\bigg{\rangle}$ (16)
+Since practically we cannot simulate $n$ to infinity, we introduce a large cut-off $M$ to truncate the series.
+For those $n<M$, after inserting some identity operators (also some special $H_{a,b}$ by using $a$ to denote identity, diagonal and off-diagonal types), we can make all operator strings have the same length $M$.
+Besides, we introduce $(M-1)$ extra sets of complete basis, and finally we have
+$Z(\beta)=\sum_{S_{M}}\sum_{\alpha}\frac{\beta^{n}(M-n)!}{M!}\prod_{p=1}^{M}\langle\alpha_{p-1}|H_{a_{p},b_{p}}|\alpha_{p}\rangle$ (17)
+where we have identify $\alpha_{M}\equiv\alpha_{0}$ and $\alpha\equiv\alpha_{0},\cdots,\alpha_{M-1}$.
+The weight in Eq.(17) is
+$W(\beta;S_{M},\alpha)=\frac{\beta^{n}(M-n)!}{M!}\prod_{p=1}^{M}\langle\alpha_{p-1}|H_{a_{p},b_{p}}|\alpha_{p}\rangle$ (18)
+and the classical configurations are given by both the states $\alpha$ and the operator string $S_{M}$.
+By replacing the weights $W(\beta^{\prime};C)$ and $W(\beta^{\prime\prime};C)$ in Eq.
+(12) with $W(\beta^{\prime};S_{M},\alpha)$ and $W(\beta^{\prime\prime};S_{M},\alpha)$, we have
+$\frac{Z(\beta^{\prime})}{Z(\beta^{\prime\prime})}=\bigg{\langle}\bigg{(}\frac{\beta^{\prime}}{\beta^{\prime\prime}}\bigg{)}^{n}\bigg{\rangle}_{\beta^{\prime\prime}}$ (19)
+for $n$ to denote the number of non-identity operators.
+Similar discussions can be applied to the quantum case.
+For $H(s)=sH_{0}+(1-s)H_{1}=H_{1}+s(H_{0}-H_{1})$, we usually have $Z(s^{\prime})/Z(s^{\prime\prime})=\langle(s^{\prime}/s^{\prime\prime})^{n_{s}}\rangle_{s^{\prime\prime}}$ within the framework of SSE, where $n_{s}$ is the number of non-identity operators related to $(H_{0}-H_{1})$.
 For instance, if
 $\begin{split}H(s)=&s\bigg{[}\sum_{i}\mathbf{S}_{i}\cdot\mathbf{S}_{i+1}\bigg{]}+(1-s)\bigg{[}\sum_{i\in\text{even}}\mathbf{S}_{i}\cdot\mathbf{S}_{i+1}\bigg{]}\\ =&\sum_{i\in\text{even}}\mathbf{S}_{i}\cdot\mathbf{S}_{i+1}+s\bigg{[}\sum_{i\in\text{odd}}\mathbf{S}_{i}\cdot\mathbf{S}_{i+1}\bigg{]}\end{split}$ (20)
-where $H_{0}=\sum_{i}\mathbf{S}_{i}\cdot\mathbf{S}_{i+1}$ and $H_{1}=\sum_{i\in\text{even}}{\mathbf{S}}_{i}\cdot{\mathbf{S}}_{i+1}$, then
+where $H_{0}=\sum_{i}\mathbf{S}_{i}\cdot\mathbf{S}_{i+1}$ and $H_{1}=\sum_{i\in\text{even}}\vec{S}_{i}\cdot\vec{S}_{i+1}$, then we can obtain
 $\begin{split}\frac{Z(s^{\prime})}{Z(s^{\prime\prime})}=&\frac{1}{Z(s^{\prime\prime})}\text{tr}\{e^{-\beta H(s^{\prime})}\}\\ =&\frac{1}{Z(s^{\prime\prime})}\sum_{\alpha,S_{M}}\bigg{(}\frac{s^{\prime}}{2}\bigg{)}^{n_{\text{odd}}}\bigg{(}\frac{1}{2}\bigg{)}^{n_{\text{even}}}\tilde{S}_{\alpha,M}\\ =&\frac{1}{Z(s^{\prime\prime})}\sum_{\alpha,S_{M}}\bigg{(}\frac{s^{\prime}}{s^{\prime\prime}}\bigg{)}^{n_{\text{odd}}}\bigg{(}\frac{s^{\prime\prime}}{2}\bigg{)}^{n_{\text{odd}}}\bigg{(}\frac{1}{2}\bigg{)}^{n_{\text{even}}}\tilde{S}_{\alpha,M}\\ =&\bigg{\langle}\bigg{(}\frac{s^{\prime}}{s^{\prime\prime}}\bigg{)}^{n_{\text{odd}}}\bigg{\rangle}_{s^{\prime\prime}}\\ \end{split}$ (21)
-where $n_{\text{odd}}$ ($n_{\text{even}}$) denotes the amounts of non-identity operators acting on the bonds with odd (even) indices in SSE simulation, and
-$\tilde{S}_{\alpha,M}=\frac{\beta^{n}(M-n)!}{M!}\prod_{i=1}^{M}\langle\alpha_{i-1}|H_{a_{i},b_{i}}|\alpha_{i}\rangle$ (4)
-corresponds to operator string $S_{M}$ and the propagated states $\{\alpha_{0},\alpha_{1},\dots\alpha_{M}\}$ along the axis of imaginary time.
-A.2 (II) Technical details of the annealing process
-For the thermal and quantum reweight-annealing scheme with a scalar parameter $p$, our target is to compute
-$\frac{Z(p_{0})}{Z(p)}=\bigg{\langle}\frac{W(p_{0})}{W(p)}\bigg{\rangle}_{p}$ (5)
-where $p_{0}$ is the reference point in the parameter space.
-For the models we consider, $p_{0}=0$.
-Practically, we take $p_{0}$ to be a sufficiently small number (e.g.
-$10^{-8}$ in our paper) in the computer.
-The adequacy can be checked by comparing the results with that of a smaller $p_{0}$ so as to see whether the results converge.
-We first take $p=\beta$ for illustration and the case that $p=s$ is similar.
-The difference will be clarified at the end of this section.
-Without loss of generality, we assume $\beta_{0}<\beta$.
-In case that $\beta_{0}$ and $\beta$ are far away in the parameter space, to ensure importance sampling, we divide $[\beta_{0},\beta]$ into $m$ intervals $[\beta_{0},\beta_{1}]$, $[\beta_{1},\beta_{2}]$, $\cdots$, $[\beta_{m-1},\beta_{m}]$, where $\beta_{m}\equiv\beta$, i.e.
-$\frac{Z(\beta_{0})}{Z(\beta)}=\prod_{k=1}^{m}\frac{Z(\beta_{k-1})}{Z(\beta_{k})}$ (6)
-Each interval $[\beta_{k-1},\beta_{k}]$ ought to be sufficiently narrow so that the distribution at $\beta_{k}$ can well samples the configurations at $\beta_{k-1}$.
-Otherwise, the sampling process would be inefficient.
-Then the natural questions following are how to partition $[\beta_{0},\beta]$ and the how many intervals we need for Eq.
-(6)
-It can be expected that if $\beta_{k}$ and $\beta_{k-1}$ are separated on the two sides of the critical point, we may need a bulk of samples to estimate $Z(\beta_{k-1})/Z(\beta_{k})$.
-However, we point out that this actually affects little to our algorithm since numerical simulations only deal with finite system, where the criticality does not result in significant inefficiency.
-We have also verified this point in our simulations.
-For example, in the 2D spin-$1/2$ XXZ model, we implemented an equal division strategy with step size $\Delta T=0.01$.
-With $20000$ Monte Carlo samples for each bin, we were able to achieve a sufficiently smooth curve for further numerical differentiation.
-However, the strategy that we divide $[\beta_{0},\beta]$ into equal intervals or a geometric progression is not economic.
-We can expect that when $\beta\to 0$ (approaching infinite high temperature), the division can be more sparse compared to that at low temperatures.
-Therefore, an economic strategy to divide the interval is needed.
-In addition to the division problem, $Z(\beta_{0})/Z(\beta)$ could be an exponentially small number exceeding the data range of computer (e.g.
-$1.7\times 10^{-308}\sim 1.7\times 10^{308}$ for double type data), and we should store $\ln[Z(\beta_{0})/Z(\beta)]$ in the computer.
-However, simply applying the logarithm to Eq.
-(6) and converting the multiplication to addition may lead to precision loss in computer calculations if $\ln[Z(\beta_{k-1})/Z(\beta_{k})]$ and $\ln[Z(\beta_{k})/Z(\beta_{k+1})]s$ differ by several orders of magnitude.
-To address the two problems above, we present a dynamical approach here. Define $\alpha_{k}\equiv\beta_{k-1}/\beta_{k}<1$, and within the framework of SSE, we have
-$\frac{Z(\beta_{k-1})}{Z(\beta_{k})}=\langle\alpha^{n_{k}}_{k}\rangle_{\beta_{k}}$ (7)
-where $n_{k}$ is the number of all non-identity operators in the SSE simulation. Our dynamical process then starts from the last point with $k=m$, and we require
-$\frac{Z(\beta_{k-1})}{Z(\beta_{k})}=\epsilon<1,$ (8)
-where the choice of $\epsilon$ should match $N_{\text{MC}}$, the number of MC steps we use for samplings, as the error in MC simulations is proportional to $1/\sqrt{N_{\text{MC}}}$.
-Next, we set $n_{k}\approx\tilde{n}_{k}=\Lambda_{k}\beta_{k}L$ ($\Lambda_{k}\beta_{k}L^{2}$) for 1D (2D) system, and $\Lambda_{k}$ can be ascertained through the examination of result convergence.
-This comes from the fact that in SSE, the energy satisfies $E_{k}=-\langle n_{k}\rangle/\beta_{k}$.
-Practically, we set $\Lambda_{k}$ to be the same for all $k$ for convenience.
-Then $Z(\beta_{k-1})/Z(\beta_{k})$ can be obtained from the simulation from
-$\begin{split}\langle\alpha_{k}^{n_{k}}\rangle=\langle(\epsilon^{1/\tilde{n}_{k}})^{n_{k}}\rangle_{\beta_{k}}\approx\epsilon,\quad\beta_{k-1}=\alpha_{k}\beta_{k}\end{split}$ (9)
-Eq.
-(9) not only requires $Z(\beta_{k-1})/Z(\beta_{k})\approx\epsilon$ for all $k$, but determines the next step in the annealing path.
-Therefore, with this dynamical annealing strategy, the division will be more sparse when $\beta$ gets smaller.
-This strategy actually requires the number divisions of scale proportional to the system size for large systems. For the 1D system as an example,
-$\beta_{k-1}=\epsilon^{\frac{1}{\Lambda\beta_{k}L}}\beta_{k}\approx\bigg{(}1+\frac{\ln\epsilon}{\Lambda\beta_{k}L}\bigg{)}\beta_{k}$ (10)
-or
-$\beta_{k}-\beta_{k-1}=-\frac{\ln\epsilon}{\Lambda L}$ (11)
-then the number of divisions is around $(\beta-\beta_{0})\Lambda L/|\ln\epsilon|$.
-One more thing we have to pay attention is that when $\beta\to 0$, we have $\alpha_{k}\to 0$, which will also exceed the range of a number saved in the computer.
-Therefore we manually set a lower bound for $\alpha_{k}$ (e.g.
-$\min{\alpha_{k}}=\epsilon$ in our simulation).
-The existence of lower bound makes the division in the region that $\beta\to 0$) denser than that Eq.
-(9) requires.
-Since most $Z(\beta_{k-1})/Z(\beta_{k})$ still at the same order of magnitude, e.g.
-$\epsilon=10^{-2}$ used for the first two example in our simulation, and the more divisions when $\beta\to 0$ only constitutes a very small proportion, this affects little to the precision.
-In Fig.
-1, we plot the number of divisions in the T-Re-An method of first example in our paper for reference, where we take $\Lambda=L$.
-Figure 1: The number of divisions in the T-Re-An method of first example in our paper, where we take $\beta=30$ and $\Lambda=L$.
-For the quantum reweight-annealing scheme, we adopt a similar procedure as introduced above.
-For $Z(s_{k-1})/Z(s_{k})=\langle(s_{k-1}/s_{k})^{n_{k,\text{odd}}}\rangle_{s_{k}}$, we can expect that when $s\to 0$, the number of the operators of the bonds with odd indices goes to zero in Eq.
-(3).
-Therefore we can take $n_{k,\text{odd}}\approx\tilde{n}_{k,\text{odd}}=\Gamma_{k}\beta s_{k}L$ ($\Gamma_{k}\beta s_{k}L^{2}$) for 1D (2D) system, and we find this hypothetical linear relation on $s$ is convenient and has sufficiently good performance.
-The number of divisions in this case similarly scales with $\mathcal{O}(\Gamma\beta L)$ for some uniform $\Gamma_{k}=\Gamma$ if in 1D system.
-A.3 (III) Computing entanglement entropy
-One important application of our method is to calculate the entanglement entropy in conjunction with the Bisognano-Wichmann theorem, which offers an excellently approximated functional form to the entanglement Hamiltonian in systems exhibiting translational symmetry.
+where $n_{\text{odd}}$ ($n_{\text{even}}$) denotes the amounts of non-identity operators acting on the bonds with odd (even) indices in SSE simulations, and
+$\tilde{S}_{\alpha,M}=\frac{\beta^{n}(M-n)!}{M!}\prod_{p=1}^{M}\langle\alpha_{p-1}|H_{a_{p},b_{p}}|\alpha_{p}\rangle.$ (22)
+Appendix B Computing entanglement entropy
+One important application of our method is to calculate the entanglement entropy in conjunction with the Bisognano-Wichmann theorem, which offers an excellently approximated functional form to the entanglement Hamiltonian in systems exhibiting translational symmetry [72, 73].
 We briefly introduce the functional form of 1D/2D system with periodic boundary condition here.
@@
 The spacing of lattice is taken to be $1$.
+Figure 7: For the 1D TFIM model at the critical point under PBC, the Von Neumann entropy obtained via calculating the thermal entropy of the corresponding BW-TFIM model at $\beta=1$.
+The central charge extracted from our Re-An simulations is $c=0.501(4)$, which is consistent with the theoretical result $0.5$.
 As it shown in Fig. 6, for an 1D/2D system under half bipartition (cornerless) with $A$ and $B$ part, if the Hamiltonian takes the form
@@
 Since it requires the value of $\ln\tilde{Z}_{A}$, our method can thus be used to calculate the Von Neumann entropy of lattice systems with translational symmetry.
-In addition to the LBW-HM model shown in the main body of the paper, we present another example of the 1D transverse field Ising model (TFIM), and calculate its Von Neumann entropy at the critical point with the reweight-annealing method.
+In addition to the LBW-HM model shown in the main body of the paper, we present an additional example of the 1D transverse field Ising model (TFIM), and calculate its Von Neumann entropy at the critical point with the reweight-annealing method.
 The exact result is $S(L)\propto(c/3)\ln L$, where $c=0.5$ is the central charge, and our method gives $c\approx 0.501(4)$, as it shown in Fig.
-3.
-Figure 3: For the 1D TFIM model at the critical point under PBC, the Von Neumann entropy obtained via calculating the thermal entropy of the corresponding BW-TFIM model at $\beta=1$.
-The central charge extracted from our Re-An simulations is $c=0.501(4)$, which is consistent with the theoretical result $0.5$.
+7.